### v1

Abstract
This paper presents and investigates an inexact proximal gradient method for solving composite convex optimization problems characterized by an objective function composed of a sum of a full domain differentiable convex function and a non-differentiable convex function. We introduce an explicit linesearch strategy that requires only a relative inexact solution of the proximal subproblem per iteration. We prove the convergence of the sequence generated by our scheme and establish its iteration complexity, considering both the functional values and a residual associated with first-order stationary solutions. Additionally, we provide numerical experiments to illustrate the practical efficacy of our method.
Keywords:
Linesearch Iteration complexity Nonsmooth and convex optimization problemProximal gradient method Relative error rule
MSC:
65K05 90C25 90C30
1 Introduction
In this paper, our focus is on addressing nonsmooth convex optimization problems characterized by the following formulation:
$\min\,F(x):=f(x)+g(x)\quad\text{subject to}\quad x\in\mathbbm{E},$ (1)
where $\mathbbm{E}$ represents a finite-dimensional Euclidean space. The function $g:\mathbbm{E}\to\overline{\mathbbm{R}}:=\mathbbm{R}\cup\{+\infty\}$ is a nonsmooth, proper, and lower semicontinuous convex function. The function $f:\mathbbm{E}\to\mathbbm{R}$ is a continuously differentiable and convex function. Throughout the paper, we denote the optimal value and solution set of problem (1) by $F_{*}$ and $S_{*}$, respectively. From now on, we assume that $S_{*}\neq\emptyset$.
Proximal gradient methods effectively solve optimization problems such as in (1). The main step in the proximal gradient method involves evaluating the proximal operator ${\bf prox}_{g}:\mathbbm{E}\to{\bf dom\,}g:=\{x\in\mathbbm{E}\mid g(x)<+\infty\}$ defined by:
${\bf prox}_{g}(y):=\underset{x\in\mathbbm{E}}{\arg\min}\;\left\{g(x)+\frac{1}{2}\left\|x-y\right\|^{2}\right\},$ (2)
[…]
Proximal Gradient Method (PGM) Let $x_{0}\in{\bf dom\,}g$. Compute $\gamma_{k}>0$ and $\bar{x}_{k}:={\bf prox}_{\gamma_{k}g}\left(x_{k}-\gamma_{k}\nabla f(x_{k})\right).$ (3) Choose $\beta_{k}\in(0,1]$ and compute $x_{k+1}:=x_{k}+\beta_{k}(\bar{x}_{k}-x_{k}).$ (4)
The coefficients $\gamma_{k}$ and $\beta_{k}$, referred to as stepsizes, can be determined based on backtracking linesearch procedures. Such strategies are essential whenever the global $L$-Lipschitz continuity of $f$ fails or even when computing an acceptable upper bound for $L$ is challenging. This situation is often encountered in numerous applications; for instance, in inverse problems based on Banach norms Bredies:2009 ; Schuster:2012 or Bregman distances such as the Kullback-Leibler divergence Salzo:2014 ; Bonettini:2012 ; BelloCruz:2016a . Moreover, even when $L$ is known, linesearches may allow for longer steps toward the solution by using local information at every iteration.
There are several possible choices for these stepsizes, each impacting the algorithm’s performance in different ways; see, for instance, Salzo:2017a ; BelloCruz:2016a ; BelloCruz:2016c ; Bello-Cruz:2021 . It is important to note that in order to compute the stepsize $\gamma_{k}$ using a backtracking linesearch at each iteration $k$, the proximal operator may need to be evaluated multiple times within the procedure. Conversely, the stepsize $\beta_{k}$ can be selected by evaluating the proximal operator only once per iteration. In this context, we will refer to explicit linesearch to describe a backtracking procedure that determines $\beta_{k}$ after setting $\gamma_{k}$ as a constant for all $k$. These explicit strategies are particularly advantageous, especially in cases where evaluating the proximal operator is challenging. The function $g$ is often complex enough that the corresponding proximal operator lacks an analytical solution. In such cases, an ad-hoc algorithm should be employed to evaluate the proximal operator inexactly. For instance, consider $\mathbb{E}=\mathbbm{R}^{n}$, and let $g:\mathbbm{R}^{n}\to\mathbbm{R}$ be defined as
$g(x)=\|x\|_{1}+\lambda\sum_{i<j}\max\{|x_{i}|,|x_{j}|\},$
with $\lambda>0$, a form encountered in sparse regression problems with grouping, as discussed in Zhong:2012 . Similarly, in the context of matrix factorization, consider $\mathbb{E}=\mathbbm{R}^{n\times m}$ for the CUR-like factorization optimization problem Barre:2022 ; JMLR:v12:mairal11a , where the goal is to approximate a matrix $W\in\mathbbm{R}^{m\times n}$ with $X\in\mathbbm{R}^{n\times m}$ having sparse rows and columns. In this case, $g:\mathbbm{R}^{n\times m}\to\mathbbm{R}$ is given by
$g(X)=\lambda_{\text{row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+\lambda_{\text{col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2},$
which will be considered in the numerical illustrations of this paper. For further examples and discussions, see BelloCruz:2017a ; Bello-Cruz:2023 ; Salzo:2012 ; Villa:2013 ; Schmidt:2011 ; Jiang:2012 ; Millan:2019 ; Barre:2022 . Of course, there are rare cases when the exact analytical solution of the proximal operator is available, such as when $g:\mathbbm{R}^{n}\to\mathbbm{R}$ as $g(x)=\|x\|_{1}$ or the indicator function of a simple convex and closed set, see, for instance, Beck:2009 ; Combettes:2005 .
Consequently, in practice, the evaluation of the proximal operator is often done inexactly. A basic inexactness criterion is to approximately evaluate the proximal operator using an exogenous sequence of error tolerance which, in general, must be summable in order to guarantee the convergence of the proximal iterations. Such a diminishing sequence is a priori chosen without using any information that may be available along with the iterations. Usually, the restrictive summability condition forces the solution of the proximal subproblem to be increasingly accurate, often more than necessary; see, for instance, Jiang:2012 ; Salzo:2012 ; Aujol:2015 ; Villa:2013 ; Schmidt:2011 . In the past two decades, relative error criteria have been considered as an effective and practical way of controlling the inexactness in solving proximal subproblems of several algorithms, including FISTA, ADMM, augmented Lagrangian, and Douglas-Rachford methods; refer to Eckstein:2013 ; Eckstein:2018 ; Alves:2020 ; Adona:2019 ; Adona:2020 for examples. Relative error criteria are often easy to verify in practice and has the advantage of exploring the available information at a specific iteration, being, therefore, an interesting alternative to the aforementioned exogenous sequences.
In this paper, we propose and analyze an inexact proximal gradient method for solving problem (1). We propose a novel relative inexactness criterion for solving the proximal subproblems which in some sense ressembles the ideas of relative error criteria introduced in Monteiro:2010 ; Solodov:1999a , but adds some new elements in order to control the objective function on the inexact solution. The proposed scheme requires only one inexact solution of the proximal subproblem per iteration, and the stepsizes are computed through a relaxed explicit linesearch procedure that takes into account the residuals obtained in the proximal subproblem and enables the iteration to address non-Lipschitz optimization problems effectively. We show that the sequence generated by our method converges to a solution of problem (1). Moreover, we establish its iteration-complexity in terms of both the function values and the residues associated to an approximate stationary solution. We also present some numerical experiments to illustrate the performance of the proposed method.
The paper is structured as follows: Section 2 presents definitions, basic facts, and auxiliary results. In Section 3, we introduce the inexact proximal gradient method with an explicit linesearch (IPG-ELS) and establish some of its fundamental properties. Section 4 establishes the iteration-complexity bounds of the IPG-ELS method in terms of functional values and a residual associated with the stationary condition for problem (1). Additionally, the full convergence of the sequence generated by the IPG-ELS method is discussed in this section. Some numerical experiments illustrating the behavior of the proposed scheme are reported in Section 5. Finally, concluding remarks are provided in Section 6.
2 Preliminary
In this section, we present some preliminary material and notations that will be used throughout this paper.
Let $g:\mathbbm{E}\to\overline{\mathbbm{R}}$ be a proper, lower semicontinuous, and convex function. For a given $\varepsilon\geq 0$, the $\varepsilon$-subdifferential of $g$ at $x\in{\bf dom\,}g=\{x\in\mathbbm{E}\mid g(x)<+\infty\}$, denoted by $\partial_{\varepsilon}g(x)$, is defined as
$\partial_{\varepsilon}g(x):=\{v\in\mathbbm{E}\mid g(y)\geq g(x)+\langle v,y-x\rangle-\varepsilon,\ \forall y\in\mathbbm{E}\}.$ (5)
When $x\notin{\bf dom\,}g$, we define $\partial_{\varepsilon}g(x)=\emptyset$. Any element $v\in\partial_{\varepsilon}g(x)$ is called an $\varepsilon$-subgradient. If $\varepsilon=0$, then $\partial_{0}g(x)$ becomes the classical subdifferential of $g$ at $x$, denoted by $\partial g(x)$. It follows immediately from (5) that
$v\in\partial_{\varepsilon}g(y),\;u\in\partial g(x)\quad\text{implies}\quad\langle v-u,y-x\rangle\geq-\varepsilon.$ (6)
We present two useful auxiliary results for $\partial_{\varepsilon}g$. The first one is the closedness of the graph of $\partial_{\varepsilon}g$ and the second is the so-called transportation formula; see Propositions 4.1.1 and 4.2.2 of Hiriart-Urruty:1993 .
Proposition 1 (Closed Graph Property)
Let $(\varepsilon_{k},x_{k},v_{k})_{k\in\mathbbm{N}}\subseteq\mathbbm{R}_{+}\times\mathbbm{E}\times\mathbbm{E}$ be a sequence converging to $(\varepsilon,x,v)$ with $v_{k}\in\partial_{\varepsilon_{k}}g(x_{k})$ for all $k\in\mathbbm{N}$. Then, $v\in\partial_{\varepsilon}g(x)$.
Proposition 2 (Transportation Formula)
With $x$ and $y$ in ${\bf dom\,}g,$ let $v\in\partial g(y)$. Then $v\in\partial g_{\epsilon}(x)$ where $\epsilon=g(x)-g(y)-\langle{v},{x-y}\rangle\geq 0$.
We now introduce a concept of approximate stationary solution to problem (1). First, note that $\bar{x}$ is a solution to problem (1) if and only if $0\in\nabla f(\bar{x})+\partial g(\bar{x})$. The concept of approximate stationary solution relaxes this inclusion by introducing a residual $v$ and enlarging $\partial g$ using $\partial_{\varepsilon}g$.
Definition 1 ($\eta$-Approximate Stationary Solution)
Given $\eta>0$, an element $\tilde{x}\in{\bf dom\,}g$ is said to be an $\eta$-approximate stationary solution to problem (1) with a residual pair $(v,\varepsilon)$ if
$v\in\nabla f(\tilde{x})+\partial_{\varepsilon}g(\tilde{x}),\quad\max\{\|v\|,\varepsilon\}\leq\eta.$ (7)
Next, we recall the definition of quasi-Fejér convergence, which is an important and well-known tool for establishing full convergence of gradient and proximal point type methods.
Definition 2 (Quasi-Fejér Convergence)
Let $S$ be a nonempty subset of $\mathbbm{E}$. A sequence $(x_{k})_{k\in\mathbbm{N}}\subseteq\mathbbm{E}$ is said to be quasi-Fejér convergent to $S$ if and only if, for every $x\in S$, there exists a summable sequence $(\delta_{k})_{k\in\mathbbm{N}}\subseteq\mathbbm{R}_{+}$ such that
$\|x_{k+1}-x\|^{2}\leq\|x_{k}-x\|^{2}+\delta_{k},\qquad\forall k\in\mathbbm{N}.$ (8)
The following result presents the main properties of quasi-Fejér convergent sequences; see Theorem $2.6$ of Bauschke:1996 .
Lemma 1 (Quasi-Féjer Convergence Properties)
If $(x_{k})_{k\in\mathbbm{N}}$ is quasi-Fejér convergent to $S$, then the following statements hold:
The sequence $(x_{k})_{k\in\mathbbm{N}}$ is bounded;
If a cluster point $\bar{x}$ of $(x_{k})_{k\in\mathbbm{N}}$ belongs to $S$, then $(x_{k})_{k\in\mathbbm{N}}$ is convergent to $\bar{x}$.
We conclude the section with a basic inequality that will be used in the next sections.
Lemma 2
For any $x,y\in\mathbbm{E}$, we have $\|x+y\|^{2}\leq 2\|x\|^{2}+2\|y\|^{2}$.
3 Inexact Proximal Gradient Method
In this section, we introduce the inexact proximal gradient method with an explicit linesearch and establish some basic properties.
Recall that, given $x_{k}\in\mathbbm{E}$, the exact solution of the proximal gradient subproblem (3) with $\gamma_{k}=1$ consists of finding $\bar{x}_{k}$ such that
$\bar{x}_{k}={\rm arg}\!\min_{x\in\mathbbm{E}}\left\{\left\langle\nabla f(x_{k}),x-x_{k}\right\rangle+g(x)+\frac{1}{2}\|x-x_{k}\|^{2}\right\}.$
This proximal subproblem is equivalent to the following monotone inclusion problem:
$0\in\nabla f(x_{k})+\partial g(\bar{x}_{k})+\bar{x}_{k}-x_{k}.$ (9)
The $k$-th iteration of the inexact proximal gradient method, which incorporates an explicit linesearch as proposed below, begins by first relaxing the above inclusion. This is achieved by finding an approximate solution $\tilde{x}_{k}$ along with a residual pair $(v_{k},\varepsilon_{k})$ satisfying
$v_{k}\in\nabla f(x_{k})+\partial_{\varepsilon_{k}}g(\tilde{x}_{k})+\tilde{x}_{k}-x_{k},$
and adheres to a specific mixed relative error criterion. This criterion is designed to ensure that the residual pair remains small, while also controlling the functional value of $g$ and the angle between $\nabla f(x_{k})$ and $v_{k}$, in addition to guiding a search direction that involves the intermediate iteration $\tilde{x}_{k}$ and the residual $v_{k}$. If the stopping criterion $\tilde{x}_{k}=x_{k}$ is not satisfied, we use a novel linesearch procedure, designed with the triple $(\tilde{x}_{k},v_{k},\varepsilon_{k})$ in mind, to find a suitable stepsize $\beta_{k}$. This stepsize is used to move from the current point $x_{k}$ towards $y_{k}-x_{k}$ where $y_{k}=\tilde{x}_{k}-v_{k}$, genegenerating the next point $x_{k+1}$.
In the following, we formally describe the proposed method, which combines a relative inexact computation of the proximal subproblem with an explicit linesearch.
[…]
Remark 1 (The Inexact Proximal Subproblem)
Note that the inexact proximal gradient subproblem (10) is equivalent to the exact proximal gradient subproblem (9) whenever $v_{k}=0$ and $\varepsilon_{k}=0$. Moreover, it is possible to always set $v_{k}$ to 00 and still introduce a novel scheme that diverges from the previous one, which was based on inexact summable error tolerance. Indeed, under this assumption, the inexact proximal subproblem (10)-(11) can be rewritten as:
$0\in\nabla f(x_{k})+\partial_{\varepsilon_{k}}g(\tilde{x}_{k})+\tilde{x}_{k}-x_{k},$ (13)
$\varepsilon_{k}\leq\frac{(1-\tau-\alpha)}{2(1+\gamma_{2})}\|x_{k}-\tilde{x}_{k}\|^{2}.$ (14)
It is worth noting that for the above subproblem, no summability assumption is made on the tolerance sequence $(\varepsilon_{k})_{k\in\mathbbm{N}}$ to achieve the convergence and complexity of the proposed method, as will be shown later. Furthermore, inclusion (10) and inequality (11) guarantee that $\tilde{x}_{k}$ and $\tilde{x}_{k}-v_{k}$ are in the ${\bf dom\,}g$. Additionally, if $g$ is the indicator function of a nonempty closed convex set $C$, then (13)–(14) can be fulfilled by finding a point $\tilde{x}_{k}\in C$ such that
$\langle x_{k}-\nabla f(x_{k})-\tilde{x}_{k},y-\tilde{x}_{k}\rangle\leq\varepsilon_{k}:=\frac{(1-\tau-\alpha)}{2(1+\gamma_{2})}\|x_{k}-\tilde{x}_{k}\|^{2},\quad\forall y\in C.$ (15)
For example, if $C$ is bounded, one can use the conditional gradient (CondG) method, a.k.a. Frank-Wolfe method Frank:1956 ; Jaggi:2013 , to obtain such point $\tilde{x}_{k}\in C$. Indeed, given $z_{t}\in C$, the $t$-th step of the CondG method, applied to solve the projection problem $\min_{y\in C}\|y-x_{k}+\nabla f(x_{k})\|^{2}/2$, first finds $w_{t}$ as a minimum of the linear function $\langle{z_{t}-x_{k}+\nabla f(x_{k})},{\cdot-z_{t}}\rangle$ over $C$ and then sets $z_{t+1}=(1-\alpha_{t})z_{t}+\alpha_{t}w_{t}$ for some $\alpha_{t}\in(0,1)$. If at an iteration $\ell$, the computed point $z_{\ell}$ satisfies the stopping criterion $\langle{z_{\ell}-x_{k}+\nabla f(x_{k})},{w_{\ell}-z_{\ell}}\rangle\geq-\varepsilon_{k}$ for the CondG method, then $\tilde{x}_{k}:=z_{\ell}$ and (15) will hold.
Remark 2 (The Explicit Lineasearch)
Note that the explicit linesearch proposed in Step 4 of the IPG-ELS is related to the one proposed in BelloCruz:2016a when $v_{k}=0$ and $\varepsilon_{k}=0$. Moreover, note that the linesearch does not evaluate the inexact proximal subproblem inside of the inner loop of Step 4.
We introduce a useful lemma that plays a crucial role in analyzing the stopping criterion of the IPG-ELS method.
Lemma 3 (Iteration Inequality Condition)
The following inequality holds for every iteration $k$.
$\left\langle\tilde{x}_{k}-x_{k},v_{k}\right\rangle+\frac{(\gamma_{1}-1)}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}\leq\frac{(1-\tau-\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2}.$ (16)
Proof It follows from (10) that
$v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k})\in\partial_{\varepsilon_{k}}g(\tilde{x}_{k}).$
Now using (5) together with the fact that $y_{k}=\tilde{x}_{k}-v_{k}$, we have
$g(y_{k})-g(\tilde{x}_{k})$ $\geq\left\langle v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k}),y_{k}-\tilde{x}_{k}\right\rangle-\varepsilon_{k}$
$=\left\langle v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k}),-v_{k}\right\rangle-\varepsilon_{k},$
which is equivalent to
$\Gamma_{k}:=g(y_{k})-g(\tilde{x}_{k})-\langle\nabla f({x}_{k}),v_{k}\rangle$ $\geq-\|v_{k}\|^{2}+\langle\tilde{x}_{k}-x_{k},v_{k}\rangle-\varepsilon_{k}.$ (17)
On the other hand, considering the definitions of $\Gamma_{k}$ and $y_{k}$, we observe that (11) is equivalent to
$\Gamma_{k}+\frac{(1+\gamma_{1})}{2}\|v_{k}\|^{2}+(1+\gamma_{2})\varepsilon_{k}\leq\frac{(1-\tau-\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2},$
which, combined with (17), yields the desired inequality (16). ∎
The following result demonstrates that the termination criterion of the IPG-ELS method, as specified in Step 3, is satisfied only when a solution to problem (1) is identified.
Lemma 4 (Stopping at a Solution)
The IPG-ELS method terminates at the $k$-th iteration if and only if $x_{k}$ is a solution to problem (1).
Proof Assume that the IPG-ELS method stops at the $k$-th iteration. In view of Step 3, we have $\tilde{x}_{k}=x_{k}$. Hence, it follows from (16) that
$\frac{(\gamma_{1}-1)}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}\leq 0.$
Since $\gamma_{1}>1$ and $\gamma_{2}\geq 1$, we obtain $v_{k}=0$ and $\varepsilon_{k}=0$. Hence, in view of (10), we get that $0\in\nabla f(x_{k})+\partial g(x_{k}),$ concluding that $x_{k}$ is a solution of problem (1).
Assume now that $x_{k}$ is a solution of problem (1). Thus, $-\nabla f(x_{k})\in\partial g(x_{k}).$ It follows from (10) that $v_{k}+x_{k}-\tilde{x}_{k}-\nabla f(x_{k})\in\partial_{\varepsilon_{k}}g(\tilde{x}_{k})$. So, the $\varepsilon$-monotonicity of $\partial_{\varepsilon}g$ in (6) implies
$\langle v_{k}+x_{k}-\tilde{x}_{k},\tilde{x}_{k}-x_{k}\rangle\geq-\varepsilon_{k},$
which is equivalent to
$\langle v_{k},\tilde{x}_{k}-x_{k}\rangle\geq\|x_{k}-\tilde{x}_{k}\|^{2}-\varepsilon_{k}.$ (18)
Hence, it follows from (16) that
$\frac{(1+\tau+\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{(\gamma_{1}-1)}{2}\|v_{k}\|^{2}+(\gamma_{2}-1)\varepsilon_{k}\leq 0.$
Since $\gamma_{1}>1$, $\gamma_{2}\geq 1$, $\alpha\geq 0$, and $\tau>0$, we conclude that $x_{k}=\tilde{x}_{k}$ and $v_{k}=0$. Hence, from (11), we also have $\varepsilon_{k}=0$. Therefore, the IPG-ELS method stops at the $k$-th iteration. ∎
In the following, we establish the well-definedness of the linesearch procedure in Step 4 of the IPG-ELS method.
Lemma 5 (Finite Linesearch Termination)
The linesearch procedure in Step 4 stops after a finite number of steps.
Proof In view of Step 3, we have $x_{k}\neq\tilde{x}_{k}$. Assume, for the sake of contradiction, that the linesearch procedure does not terminate after a finite number of steps. Thus, for all $\beta\in\{1,\theta,\theta^{2},\ldots\}$, we have
$f\left(x_{k}+\beta(y_{k}-x_{k})\right)>f(x_{k})+\beta\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{\beta\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\beta\|v_{k}\|^{2}+\beta\gamma_{2}\varepsilon_{k},$
or, equivalently,
$\frac{f\left(x_{k}+\beta(y_{k}-x_{k})\right)-f(x_{k})}{\beta}-\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle>\frac{\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}.$
Given that $f$ is differentiable, the left-hand side of the above inequality approaches zero as $\beta\downarrow 0$, leading to the conclusion that
$0\geq\frac{\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}.$
This implies $x_{k}=\tilde{x}_{k}$ contradicting the assumption that $x_{k}\neq\tilde{x}_{k}$. ∎
The subsequent analysis investigates the linesearch procedure introduced in Step 4 of the IPG-ELS method, assuming that the gradient of the function $f$, $\nabla f$, is Lipschitz continuous. We aim to establish an upper bound for the number of iterations required by the linesearch procedure in Step 4.
Lemma 6 (Lipschitz Condition and Linesearch Complexity)
Assume that $f$ has an $L$-Lipschitz continuous gradient and that $x_{k}$ is not a solution to problem (1). Then, any $\beta\leq\tau/(2L)$ satisfies (12). As a consequence, the linesearch procedure in Step 4 of the IPG-ELS method stops in at most
$\ell:=\left\lceil\frac{\ln(\min\{\tau/(2L),1\})}{\ln(\theta)}\right\rceil$ (19)
iterations.
Proof Since $\nabla f$ is $L$-Lipschitz continuous, for any $\beta>0$, we have
$f(x_{k}+\beta(y_{k}-x_{k}))-f(x_{k})-\beta\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle$ $\leq\frac{L\beta^{2}}{2}\|y_{k}-x_{k}\|^{2}.$
Hence, if $\beta\leq\tau/(2L)$, we conclude that
$f(x_{k}+\beta(y_{k}-x_{k}))-f(x_{k})-\beta\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle$ $\leq\frac{\tau\beta}{4}\|y_{k}-x_{k}\|^{2}=\frac{\tau\beta}{4}\|x_{k}-\tilde{x}_{k}+v_{k}\|^{2}$
$\leq\frac{\beta\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\tau}{2}\beta\|v_{k}\|^{2},$
using Lemma 2 in the last inequality. Since $\gamma_{1}>1>\tau$ and $\gamma_{2}\geq 0$, we have that (12) holds, thereby proving the first statement of the lemma. The last statement follows from the first one, given that the natural number $\ell$, defined in (19), satisfies $\beta_{\ell}:=\theta^{\ell}\leq\min\{\tau/(2L),1\}$. ∎
This lemma provides a sufficient condition to ensure that the lower bound of the sequence generated by our linesearch is greater than 00. Specifically, if $\nabla f$ is $L$-Lipschitz, then the step sizes $\beta_{k}$, produced through the line search (12), are guaranteed to be bounded below by a positive constant $\beta>0$, i.e., $\beta_{k}\geq\beta>0$ for all $k\in\mathbb{N}$. Moreover, it is possible to relax the global Lipschitz condition to something local, such as $\nabla f$ being locally Lipschitz continuous around any solution of problem (1), as was done in Proposition 5.4(ii) of BelloCruz:2016a .
4 Convergence and Complexity Analyses of the IPG-ELS Method
In this section, we focus on analyzing the convergence properties of the IPG-ELS method. We establish the convergence and its iteration complexity in terms of functional values and a residual associated with an approximated solution, as defined in Definition 1.
We begin this section by presenting a result that is fundamental for establishing the convergence and the iteration complexity of the IPG-ELS method.
Proposition 3 (Key Inequality for the IPG-ELS Method)
For every $x\in{\bf dom\,}\,g$ and $k\in\mathbbm{N}$, we have
$2\beta_{k}[F(x_{k})-F(x)]+2({F(x_{k+1})-F(x_{k})})\leq\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}-{\alpha\beta_{k}}\|x_{k}-\tilde{x}_{k}\|^{2}.$ (20)
Additionally, the functional value sequence $(F(x_{k}))_{k\in\mathbbm{N}}$ is decreasing and convergent, and $\sum_{k=0}^{\infty}\|x^{k}-x^{k+1}\|^{2}<+\infty$.
Proof Let $x\in{\bf dom\,}\,g$ and $k\in\mathbb{N}$. In view of the inexact relative inclusion (10), we have $v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k})\in\partial_{\varepsilon_{k}}g(\tilde{x}_{k})$, and hence the definition of $\partial_{\varepsilon}g$ in (5) implies that
$g(x)\geq g(\tilde{x}_{k})+\langle v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k}),x-\tilde{x}_{k}\rangle-\varepsilon_{k}.$ (21)
Since $f$ is convex, we have $f(x)-f(x_{k})\geq\langle\nabla f(x_{k}),x-x_{k}\rangle$. Adding the above two inequalities, using $f+g=F$, and simplifying the resulting expression, we obtain
$F(x)-F(x_{k})\geq g(\tilde{x}_{k})-g(x_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-x_{k}\rangle-\langle{{x}_{k}-\tilde{x}_{k}},{\tilde{x}_{k}-x}\rangle+\langle{v_{k}},{x-\tilde{x}_{k}}\rangle-\varepsilon_{k}.$
Combining the above inequality with the identity
$-\langle{{x}_{k}-\tilde{x}_{k}},{\tilde{x}_{k}-x}\rangle=\frac{1}{2}\left[\|x_{k}-\tilde{x}_{k}\|^{2}+\|\tilde{x}_{k}-x\|^{2}-\|x_{k}-x\|^{2}\right],$
we have
$F(x)-F(x_{k})\geq$ $\;g(\tilde{x}_{k})-g(x_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-x_{k}\rangle+\frac{1}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+\frac{1}{2}\left[\|\tilde{x}_{k}-x\|^{2}-\|x_{k}-x\|^{2}\right]+\langle{v_{k}},{x-\tilde{x}_{k}}\rangle-\varepsilon_{k},$
or, equivalently,
$F(x)-F(x_{k})+\frac{F(x_{k})-F(x_{k+1})}{\beta_{k}}\geq$ $\frac{g(x_{k})-g(x_{k+1})}{\beta_{k}}+g(\tilde{x}_{k})-g(x_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-y_{k}\rangle$
$+\frac{f(x_{k})-f(x_{k+1})}{\beta_{k}}+\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{1}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+\frac{1}{2}\left[\|x-(\tilde{x}_{k}-v_{k})\|^{2}-\|x_{k}-x\|^{2}\right]-\frac{1}{2}\|v_{k}\|^{2}-\varepsilon_{k}.$
Since $x_{k+1}=x_{k}+\beta_{k}(y_{k}-x_{k})$ and $g$ is convex, we have $g(x_{k+1})-g(x_{k})\leq\beta_{k}(g(y_{k})-g(x_{k}))$. Hence, combining the last two inequalities and the fact that $y_{k}=\tilde{x}_{k}-v_{k}$, we get
$F(x)-F(x_{k})+\frac{F(x_{k})-F(x_{k+1})}{\beta_{k}}\geq$ $\frac{1}{2}\left[\|x-y_{k}\|^{2}-\|x_{k}-x\|^{2}\right]$
$+\frac{f(x_{k})-f(x_{k+1})}{\beta_{k}}+\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{1}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+g(\tilde{x}_{k})-g(y_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-y_{k}\rangle-\frac{1}{2}\|v_{k}\|^{2}-\varepsilon_{k}.$
or, equivalently,
$F(x)-F(x_{k})$ $+\frac{F(x_{k})-F(x_{k+1})}{\beta_{k}}\geq\frac{1}{2}\left[\|x-y_{k}\|^{2}-\|x_{k}-x\|^{2}\right]+\frac{\alpha}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+\frac{f(x_{k})-f(x_{k+1})}{\beta_{k}}+\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}$
$+g(\tilde{x}_{k})-g(y_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-y_{k}\rangle-\frac{(1+\gamma_{1})}{2}\|v_{k}\|^{2}-(1+\gamma_{2})\varepsilon_{k}$
$+\frac{(1-\tau-\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2}.$
Now, using the linesearch procedure of Step 4, we obtain
$F(x)-F(x_{k})+\frac{F(x_{k})-F(x_{k+1})}{\beta_{k}}\geq$ $\frac{1}{2}\left[\|x-y_{k}\|^{2}-\|x_{k}-x\|^{2}\right]+\frac{\alpha}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+g(\tilde{x}_{k})-g(y_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-y_{k}\rangle$
$-\frac{(1+\gamma_{1})}{2}\|v_{k}\|^{2}-(1+\gamma_{2})\varepsilon_{k}+\frac{(1-\tau-\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2}.$
It follows from the above inequality and (11) that
$\beta_{k}[F(x)-F(x_{k})]+{F(x_{k})-F(x_{k+1})}\geq\frac{\beta_{k}}{2}\left[\|x-y_{k}\|^{2}-\|x_{k}-x\|^{2}\right]+\frac{\alpha\beta_{k}}{2}\|x_{k}-\tilde{x}_{k}\|^{2}.$ (22)
On the other hand, using the identity $x_{k+1}-x=(1-\beta_{k})(x_{k}-x)+\beta_{k}(y_{k}-x)$ and the strong convexity of $\|\cdot\|^{2}$, we have
$\|x_{k+1}-x\|^{2}\leq(1-\beta_{k})\|x_{k}-x\|^{2}+\beta_{k}\|y_{k}-x\|^{2}-(1-\beta_{k})\beta_{k}\|x_{k}-y_{k}\|^{2},$
which implies
$\beta_{k}\left(\|y_{k}-x\|^{2}-\|x_{k}-x\|^{2}\right)\geq\|x_{k+1}-x\|^{2}-\|x_{k}-x\|^{2}.$
Therefore, the proof of (20) follows by combining the latter inequality with (22). The last statement of the proposition follows immediately from (20) with $x=x_{k}$ that
$02({F(x_{k+1})-F(x_{k})})\leq-\|x_{k+1}-x_{k}\|^{2}<$ 0 . (23)
So, $(F(x_{k}))_{k\in\mathbbm{N}}$ is decreasing and convergent because it is bounded from below by $F(x_{*})$. Moreover, the last inequality implies that $\sum_{k=0}^{\infty}\|x^{k}-x^{k+1}\|^{2}<+\infty$. ∎
Next, we establish the convergence rate of the functional values sequence $(F(x_{k}))_{k\in\mathbbm{N}}$.
Theorem 4.1 (Convergence Rate of the IPG-ELS Method)
Let $(x_{k})_{k\in\mathbbm{N}}$ and $(\beta_{k})_{k\in\mathbbm{N}}$ be generated by the IPG-ELS method. Assume that there exists $\beta>0$ such that $\beta_{k}\geq\beta$ for all $k\in\mathbbm{N}$. Then, for all $k\in\mathbbm{N}$, we have
$F(x_{k})-F_{*}\leq\frac{{\rm dist}(x_{0},S_{*})^{2}+2\left(F(x_{0})-F_{*}\right)}{2\beta(k+1)}.$ (24)
Proof For any $\ell\in\mathbbm{N}$ and $x_{*}\in S_{*}$, it follows from Proposition 3 with $k=\ell$ and $x=x_{*}$ that
$F(x_{\ell})-F(x_{*})\leq\frac{1}{2\beta_{\ell}}\left[\|x_{\ell}-x_{*}\|^{2}-\|x_{\ell+1}-x_{*}\|^{2}+2\left(F(x_{\ell})-F(x_{\ell+1})\right)\right]$
Summing the above inequality over $\ell=0,1,\ldots,k$, we have
$\sum_{\ell=0}^{k}\left(F(x_{\ell})-F(x_{*})\right)$ $\leq\frac{1}{2\beta}\left[\|x_{0}-x_{*}\|^{2}-\|x_{k+1}-x_{*}\|^{2}+2\left(F(x_{0})-F(x_{k+1})\right)\right]$
$\leq\frac{1}{2\beta}\left[\|x_{0}-x_{*}\|^{2}+2\left(F(x_{0})-F(x_{*})\right)\right].$ (25)
Since in view of the last statement of Proposition 3, we have $F(x_{\ell+1})\leq F(x_{\ell})$ for any $\ell\in\mathbbm{N}$, it follows from (4) that
$(k+1)\left(F(x_{k})-F(x_{*})\right)\leq\frac{1}{2\beta}\left(\|x_{0}-x_{*}\|^{2}+2\left[F(x_{0})-F(x_{*})\right]\right).$
Since $x_{*}\in S_{*}$ is arbitrary, the proof of (24) follows. ∎
Remark 3 (Complexity of $\eta$-Approximate Solution)
It follows from Theorem 4.1 that every accumulation point of the sequence $(x_{k})_{k\in\mathbbm{N}}$ is a solution to problem (1). Moreover, given any $\eta>0$, in at most $k=\mathcal{O}(1/\eta^{2})$ iterations, the IPG-ELS method generates an $\eta$-approximate solution $x_{k}$ to problem (1), in the sense that $F(x_{k})-F_{*}\leq\eta$.
Theorem 4.2 (Complexity of $\eta$-Approximate Stationary Solution)
Consider $(x_{k})_{k\in\mathbbm{N}}$, $(\tilde{x}_{k})_{k\in\mathbbm{N}}$, $(v_{k})_{k\in\mathbbm{N}}$, and $(\varepsilon_{k})_{k\in\mathbbm{N}}$ generated by the IPG-ELS method and define $w_{k}:=v_{k}+x_{k}-\tilde{x}_{k}+\nabla f(\tilde{x}_{k})-\nabla f(x_{k})$, for every $k\in\mathbbm{N}$. Then, we have
$w_{k}\in\nabla f(\tilde{x}_{k})+\partial_{\varepsilon_{k}}g(\tilde{x}_{k})\subseteq\partial_{\varepsilon_{k}}F(\tilde{x}_{k}),\qquad\forall k\in\mathbbm{N}.$ (26)
Additionally, if $\alpha\in(0,1-\tau)$, and $\nabla f$ is Lipschitz continuous on ${\bf dom\,}g$, then, given a tolerance $\eta>0$, the IPG-ELS method generates an $\eta$-approximate stationary solution $\tilde{x}_{k}$ to problem (1) with residues $(w_{k},\varepsilon_{k})$, in the sense of Definition 1, in at most $k=O\left(1/\eta^{2}\right)$ iterations.
Proof The first inclusion in (26) follows immediately from (10) and the definition of $w_{k}$, whereas the second inclusion in (26) follows from the definitions of $F$ and $\partial_{\varepsilon}F$.
Now let $x_{*}$ be the projection of $x_{0}$ onto $S_{*}$ and let $d_{0}:=\|x_{0}-x_{*}\|$. As it was observed in Lemma 6, the Lipschitz continuity of $\nabla f$ implies that there exists $\beta>0$ such that $\beta_{k}\geq\beta$ for all $k$. It follows from Proposition 3 with $k=\ell\in\mathbbm{N}$ and $x=x_{*}\in S_{*}$ that
$\alpha\beta\|x_{\ell}-\tilde{x}_{\ell}\|^{2}\leq\left(\|x_{\ell}-x_{*}\|^{2}-\|x_{\ell+1}-x_{*}\|^{2}+2\left[F(x_{\ell})-F(x_{\ell+1})\right]\right)$ (27)
for all $\ell\in\mathbbm{N}$. Summing the above inequality over $\ell=0,1,\ldots,k$, and using that $x_{*}\in S_{*}$, we have
$\alpha\beta\sum_{\ell=0}^{k}\|x_{\ell}-\tilde{x}_{\ell}\|^{2}\leq\|x_{0}-x_{*}\|^{2}+2\left[F(x_{0})-F(x_{*})\right].$
Hence, since $d_{0}=\|x_{0}-x_{*}\|$, we see that there exists ${\ell_{k}}\leq k$ such that
$\|\tilde{x}_{{\ell_{k}}}-x_{{\ell_{k}}}\|^{2}\leq\frac{d_{0}^{2}+2[F(x_{0})-F(x_{*})]}{\alpha\beta(k+1)}.$ (28)
On the other hand, since $\tau,\alpha>0$, if follows from (16) that, for every $\ell\in\mathbbm{N}$,
$\frac{(\gamma_{1}-1)}{2}\|v_{\ell}\|^{2}+\gamma_{2}\varepsilon_{\ell}$ $\leq\frac{\|x_{\ell}-\tilde{x}_{\ell}\|^{2}}{2}+\langle x_{\ell}-\tilde{x}_{\ell},v_{\ell}\rangle$
$\leq\frac{\|x_{\ell}-\tilde{x}_{\ell}\|^{2}}{2}+\frac{\|x_{\ell}-\tilde{x}_{\ell}\|^{2}}{\gamma_{1}-1}+\frac{(\gamma_{1}-1)}{4}\|v_{\ell}\|^{2},$
where the last inequality is due to Cauchy-Schwarz inequality and the fact that $ab\leq sa^{2}/2+b^{2}/(2s)$ for any $a,b\in\mathbbm{R}$ and $s>0$, in particular, with $a=\|v_{\ell}\|$, $b=\|\tilde{x}_{\ell}-x_{\ell}\|$, and $s=(\gamma_{1}-1)/2$. Hence, we have
$\frac{(\gamma_{1}-1)}{4}\|v_{\ell}\|^{2}+\gamma_{2}\varepsilon_{\ell}\leq\left(\frac{\gamma_{1}+1}{2(\gamma_{1}-1)}\right)\|\tilde{x}_{\ell}-x_{\ell}\|^{2},\qquad\forall\ell\in\mathbbm{N}.$ (29)
Now, from the definition of $w_{\ell}$, (29), the Cauchy-Schwarz inequality and the Lipschitz continuity of $\nabla f$, we have, for every $\ell\in\mathbbm{N}$,
$\|w_{\ell}\|$ $\leq\|v_{\ell}\|+\|{x}_{\ell}-\tilde{x}_{\ell}\|+\|\nabla f(x_{\ell})-\nabla f(\tilde{x}_{\ell})\|$
$\leq\left[\frac{\sqrt{2(\gamma_{1}+1)}}{\gamma_{1}-1}+1+L\right]\|x_{\ell}-\tilde{x}_{\ell}\|.$ (30)
Moreover, it follows from (29) that
$\varepsilon_{\ell}\leq\left(\frac{\gamma_{1}+1}{2\gamma_{2}(\gamma_{1}-1)}\right)\|x_{\ell}-\tilde{x}_{\ell}\|^{2}.$ (31)
Hence, it follows from (28), (4), and (31) with $\ell=\ell_{k}$ and $m_{0}:=d_{0}^{2}+2(F(x_{0})-F(x_{*}))$ that
$w_{\ell_{k}}\leq\left[\frac{\sqrt{2(\gamma_{1}+1)}}{\gamma_{1}-1}+1+L\right]\frac{\sqrt{m_{0}}}{\sqrt{\alpha\beta(k+1)}},\qquad\varepsilon_{\ell_{k}}\leq\left(\frac{\gamma_{1}+1}{2\gamma_{2}(\gamma_{1}-1)}\right)\frac{m_{0}}{\alpha\beta(k+1)}$
which in turn implies that
$w_{\ell_{k}}=\mathcal{O}(1/\sqrt{k}),\quad\varepsilon_{\ell_{k}}=\mathcal{O}(1/k).$ (32)
Thus, the last statement of the theorem follows from (32) and the first inclusion in (26). ∎
We end this section by establishing the full convergence of the sequence $(x_{k})_{k\in\mathbbm{N}}$ to a solution of problem (1). The proof is based on the quasi-Fejér convergence of the sequence $(x_{k})_{k\in\mathbbm{N}}$ to the set $S_{*}$, as defined in Definition 2.
Theorem 4.3 (Convergence for the IPG-ELS Method)
The sequence $(x_{k})_{k\in\mathbbm{N}}$ generated by the IPG-ELS method converges to a point in $S_{*}$.
Proof By employing Proposition 3 at $x=x_{*}\in S_{*}\subseteq{\bf dom\,}g$, we have
$\|x_{k+1}-x_{*}\|^{2}\leq\|x_{k}-x_{*}\|^{2}+2\left[F(x_{k})-F(x_{k+1})\right]\quad\mbox{for all}\quad k\in\mathbbm{N}.$ (33)
We set $\delta_{k}:=2\left[F(x_{k})-F(x_{k+1})\right]\geq 0$, and we will prove that $(\delta_{k})_{k\in\mathbbm{N}}$ is a summable sequence. In fact,
$\sum_{k=0}^{\infty}\delta_{k}=$ $2\sum_{k=0}^{\infty}\Big{[}F(x_{k})-F(x_{k+1})\Big{]}\leq 2\Big{[}F(x_{0})-\lim_{k\to\infty}F(x_{k+1})\Big{]}$
$\leq$ $2\Big{[}F(x_{0})-F(x_{*})\Big{]}<+\infty.$
This together with (33) tells us that the sequence $(x_{k})_{k\in\mathbbm{N}}$ is quasi-Fejér convergent to $S_{*}$ via Definition 2. By Lemma 1(a), the sequence $(x_{k})_{k\in\mathbbm{N}}$ is bounded. Let $\bar{x}$ be a cluster point of $(x_{k})_{k\in\mathbbm{N}}$. Hence there exists a subsequence $(x_{\ell_{k}})_{k\in\mathbbm{N}}$ of $(x_{k})_{k\in\mathbbm{N}}$ converging to $\bar{x}$.
Now we proceed by considering the two possible cases:
Case 1. The sequence $\left(\beta_{\ell_{k}}\right)_{k\in\mathbbm{N}}$ does not converge to 00, i.e., there exist some $\beta>0$ and a subsequence of $\left(\beta_{\ell_{k}}\right)_{k\in\mathbbm{N}}$ (without relabeling) such that
$\beta_{\ell_{k}}\geq\beta,\quad\forall\,k\in\mathbbm{N}.$ (34)
By using Proposition 3 with $x=x_{*}\in S_{*}$, we get
$\beta_{k}\left[F(x_{k})-F(x_{*})\right]\leq$ $\frac{1}{2}(\|x_{k}-x_{*}\|^{2}-\|x_{k+1}-x_{*}\|^{2})+F(x_{k})-F(x_{k+1}).$
Summing for $k=0,1,\ldots,m$, the above inequality implies that
$\sum_{k=0}^{m}\beta_{k}\left[F(x_{k})-F(x_{*})\right]\leq$ $\frac{1}{2}(\|x_{0}-x_{*}\|^{2}-\|x_{m+1}-x_{*}\|^{2})+F(x_{0})-F(x_{m+1})$
$\leq$ $\frac{1}{2}\|x_{0}-x_{*}\|^{2}+F(x_{0})-F(x_{*}).$
By taking $m\to+\infty$ and using the fact that $F(x_{k})\geq F(x_{*})$ and (34), we obtain that
$\beta\sum_{k=0}^{\infty}\left[F(x_{\ell_{k}})-F(x_{*})\right]\leq\sum_{k=0}^{\infty}\beta_{\ell_{k}}\left[F(x_{\ell_{k}})-F(x_{*})\right]\leq\sum_{k=0}^{\infty}\beta_{k}\left[F(x_{k})-F(x_{*})\right]<+\infty,$
[…]
$F(x_{*})\leq F(\bar{x})\leq\liminf_{k\rightarrow\infty}F(x_{\ell_{k}})=\lim_{k\rightarrow\infty}F(x_{\ell_{k}})=F(x_{*}),$
which yields $F(\bar{x})=F(x_{*})$ and thus $\bar{x}\in S_{*}$.
[…]
$\hat{x}_{k+1}:=x_{k}+\hat{\beta}_{k}(y_{k}-x_{k})=(1-\hat{\beta}_{k})x_{k}+\hat{\beta}_{k}y_{k},$ (35)
where $y_{k}=\tilde{x}_{k}-v_{k}$. It follows from the definition of the linesearch that
$f(\hat{x}_{k+1})>f(x_{k})+\hat{\beta}_{k}\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{\tau}{2}\hat{\beta}_{k}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\hat{\beta}_{k}\|v_{k}\|^{2}+\hat{\beta}_{k}\gamma_{2}\varepsilon_{k}.$ (36)
It follows from the convexity of $f$, the fact that $\gamma_{1}>1>\tau$, and the positiveness of the term $\hat{\beta}_{k}\gamma_{2}\varepsilon_{k}$ that
$\langle\nabla f(\hat{x}_{k+1}),\hat{x}_{k+1}-x_{k}\rangle\geq f(\hat{x}_{k+1})-f(x_{k})>\hat{\beta}_{k}\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\hat{\beta}_{k}\frac{\tau}{2}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)$
which, together with (35), yields
$\hat{\beta}_{k}\frac{\tau}{2}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)$ $<\hat{\beta}_{k}\langle\nabla f(\hat{x}_{k+1})-\nabla f(x_{k}),y_{k}-x_{k}\rangle$
$\leq\hat{\beta}_{k}\|\nabla f(\hat{x}_{k+1})-\nabla f(x_{k})\|\|y_{k}-x_{k}\|$
$=\hat{\beta}_{k}\|\nabla f(\hat{x}_{k+1})-\nabla f(x_{k})\|\|x_{k}-\tilde{x}_{k}+v_{k}\|.$
Now it follows from Lemma 2 that
$\|x_{k}-\tilde{x}_{k}+v_{k}\|^{2}\leq 2\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right).$
Hence,
$\hat{\beta}_{k}\frac{\tau}{2}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)<\hat{\beta}_{k}\|\nabla f(\hat{x}_{k+1})-\nabla f(x_{k})\|\cdot\sqrt{2}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)^{\frac{1}{2}},$
which, due to the positiveness of $\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}$, yields
$\frac{\tau\sqrt{2}}{4}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)^{\frac{1}{2}}\leq\|\nabla f(\hat{x}_{k+1})-\nabla f(x_{k})\|.$ (37)
Note that $\|\hat{x}_{k+1}-x_{k}\|=\|\hat{\beta}_{k}(y_{k}-x_{k})\|=\frac{\hat{\beta}_{k}}{\beta_{k}}\|x_{k+1}-x_{k}\|=\frac{1}{\theta}\|x_{k+1}-x_{k}\|,$ which combined with the last statement of Proposition 3 give us that $\|\hat{x}_{\ell_{k}+1}-x_{\ell_{k}}\|\to 0$ as $k\to+\infty$. Since $\nabla f$ is continuous, we have $\|\nabla f(\hat{x}_{\ell_{k}+1})-\nabla f(x_{\ell_{k}})\|\to 0$ as $k\to+\infty$. From (37), it is derived that
$\lim_{k\rightarrow+\infty}\,\|x_{\ell_{k}}-\tilde{x}_{\ell_{k}}\|^{2}+\|v_{\ell_{k}}\|^{2}\leq 0,$ (38)
therefore, $\lim_{k\rightarrow+\infty}\,\|x_{\ell_{k}}-\tilde{x}_{\ell_{k}}\|=0$ and $\lim_{k\rightarrow+\infty}\|v_{\ell_{k}}\|=0$. Additionally, we can use (11) to show that $\lim_{k\rightarrow+\infty}\varepsilon_{\ell_{k}}=0$. Thus, $\bar{x}$ is also an accumulation point of the sequence $(\tilde{x}_{k})_{k\in\mathbb{N}}$, and without loss of generality, we can assume that $\tilde{x}_{\ell_{k}}\to\bar{x}$ as $k\to+\infty$. Moreover, we have that
$\lim_{k\rightarrow+\infty}\|\nabla f(x_{\ell_{k}})-\nabla f(\tilde{x}_{\ell_{k}})\|=0.$ (39)
Now, using (11), we obtain
$w_{k}:=x_{k}-\tilde{x}_{k}+v_{k}+\nabla f(\tilde{x}_{k})-\nabla f(x_{k})\in\nabla f(\tilde{x}_{k})+\partial_{\varepsilon_{k}}g(\tilde{x}_{k})\subseteq\partial_{\varepsilon_{k}}F(\tilde{x}_{k}).$
Furthermore, since $v_{\ell_{k}}$ converges to 00 as indicated by (38), and by applying the triangular inequality, we obtain
$\|w_{k}\|=\|x_{k}-\tilde{x}_{k}+v_{k}+\nabla f(\tilde{x}_{k})-\nabla f(x_{k})\|\leq\|x_{k}-\tilde{x}_{k}\|+\|v_{k}\|+\|\nabla f(\tilde{x}_{k})-\nabla f(x_{k})\|,$
which implies, via (38) and (39), that $w_{\ell_{k}}\in\partial_{\varepsilon_{\ell_{k}}}F(\tilde{x}_{\ell_{k}})$ also converges to 00. Consequently, the convergence of $\tilde{x}_{\ell_{k}}$ to $\bar{x}$ and $\varepsilon_{\ell_{k}}$ to 00, combined with the closedness of the graph of $\partial F$ in Proposition 1, gives us that $0\in\partial F(\bar{x})$. This is equivalent to stating that $\bar{x}\in S_{*}$.
In all the cases considered above, we have shown that $\bar{x}$, an accumulation point of the sequence $(x_{k})_{k\in\mathbb{N}}$, belongs to $S_{*}$. Proposition 1(b) implies that $(x_{k})_{k\in\mathbb{N}}$ converges to an optimal solution in $S_{*}$. ∎
5 Numerical Experiments
In this section, we investigate the numerical behavior of the IPG-ELS method in solving the CUR-like factorization optimization problem JMLR:v12:mairal11a . Consider $\mathbbm{E}=\mathbbm{R}^{n\times m}$. Given a matrix $W\in\mathbbm{R}^{m\times n}$, the objective is to find a sparse rows and columns matrix, $X\in\mathbbm{R}^{n\times m}$, such that $WXW$ approximates $W$. This problem can be formulated as the following splitting optimization problem:
$\min_{X\in\mathbbm{R}^{n\times m}}\frac{1}{2}\|W-WXW\|_{F}^{2}+\lambda_{\text{row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+\lambda_{\text{col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2},$
where $\|\cdot\|_{F}$ denotes the Frobenius norm, and $X^{(i)}$ and $X_{(j)}$ denote the $i$-th row and $j$-th column of $X$, respectively. This problem is a special case of problem (1) with
$f(X):=\frac{1}{2}\|W-WXW\|_{F}^{2},\quad g(X):=\lambda_{\text{row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+\lambda_{\text{col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2}.$
In this case, the gradient of $f$ is given by $\nabla f(X)=W^{T}(W-WXW)W^{T}$ and has a Lipschitz constant $L=\|W^{T}W\|^{2}$. However, the proximal operator of $g$ does not have a closed-form solution. Following the approach in Schmidt:2011 (see also Barre:2022 ; Zhou:2022 ), a triple $(\tilde{X}_{k},V_{k},\varepsilon_{k})$ satisfying (10)-(11) can be obtained using the Dykstra-like algorithm (Bauschke:2008c, , Theorem 3.3). This algorithm is applied to the proximal subproblem.
$\min_{X\in\mathbbm{R}^{n\times m}}\frac{1}{2}\|X-Z\|^{2}_{F}+{\lambda_{\rm row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+{\lambda_{\rm col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2},$ (40)
where $Z:=X_{k}-\nabla f(X_{k}),$ with the initial points $Z_{0}=Z,\quad P_{0}=0,\quad Q_{0}=0$, generates the sequences
$\left\{\begin{array}[]{l}Y_{\ell}={\bf prox}_{g_{1}}(Z_{\ell}+P_{\ell})\\ P_{\ell+1}=Z_{\ell}+P_{\ell}-Y_{\ell}\end{array}\right.\quad\mbox{and}\quad\left\{\begin{array}[]{l}Z_{\ell+1}={\bf prox}_{g_{2}}(Y_{\ell}+Q_{\ell})\\ Q_{\ell+1}=Y_{\ell}+Q_{\ell}-Z_{\ell+1},\end{array}\right.$ (41)
where $g_{1}(X)=\lambda_{\rm col}\sum_{j=1}^{m}\|X_{(j)}\|_{2}$ and $g_{2}(X)=\lambda_{\rm row}\sum_{i=1}^{n}\|X^{(i)}\|_{2}$. Recall that:
${\bf prox}_{g_{1}}(X)_{(j)}=\max\left\{1-\frac{\lambda_{\text{col}}}{\|X_{(j)}\|_{2}},0\right\}X_{(j)},\;\text{for }\quad j=1,\ldots,m$
and
${\bf prox}_{g_{2}}(X)^{(i)}=\max\left\{1-\frac{\lambda_{\text{row}}}{\|X^{(i)}\|_{2}},0\right\}X^{(i)},\;\text{for }\quad i=1,\ldots,n.$
Using now the definition of $Y_{\ell}$, as presented in (41), we have that $Z_{\ell}+P_{\ell}-Y_{\ell}\in\partial g_{1}(Y_{\ell})$. This observation, combined with Proposition 2, yields
$Z_{\ell}+P_{\ell}-Y_{\ell}\in\partial_{\epsilon_{\ell}}g_{1}(Z_{\ell+1}),\quad\mbox{where}\quad\epsilon_{\ell}=g_{1}(Z_{\ell+1})-g_{1}(Y_{\ell})-\langle{Z_{\ell}+P_{\ell}-Y_{\ell}},{Z_{\ell+1}-Y_{\ell}}\rangle\geq 0.$
On the other hand, it follows from the definition of $Z_{\ell+1}$ in (41) that
$Y_{\ell}+Q_{\ell}-Z_{\ell+1}\in\partial g_{2}(Z_{\ell+1}).$
Thus, it follows from the last two inclusions and the definition of $Z$ that
$V_{\ell}:=Z_{\ell}+P_{\ell}+Q_{\ell}-Z\in\partial_{\epsilon_{\ell}}g(Z_{\ell+1})+Z_{\ell+1}-X_{\ell}+\nabla f(X_{\ell}).$
Therefore, if the Dykstra-like algorithm is stopped when
$g(Z_{\ell}-V_{\ell})-g(Z_{\ell})-\langle\nabla f(X_{\ell}),V_{\ell}\rangle+\frac{(1+\gamma_{1})}{2}\|V_{\ell}\|^{2}+(1+\gamma_{2})\varepsilon_{\ell}\leq\frac{(1-\tau-\alpha)}{2}\|X_{\ell}-Z_{\ell}\|^{2},$
then the triple $(\tilde{X}_{k},V_{k},\varepsilon_{k}):=(Z_{\ell},V_{\ell},\epsilon_{\ell})$ satisfies (10)-(11).
Considering that the IPG-ELS method integrates two effective strategies: (i) permitting the subproblem to be solved inexactly to meet a relative approximation criterion, and (ii) employing an explicit linesearch procedure that computes the proximal operator only once per iteration, our primary goal is to demonstrate that, in certain cases, the proposed method surpasses the proximal gradient method that employs only one of these strategies. Consequently, we compare the new algorithm with two alternative schemes: an exact version of the IPG-ELS method, denoted by PG-ELS, which corresponds to IPG-ELS with $\gamma_{1}=\gamma_{2}=0$, $\theta=0.5$, $\tau=1$, and $\max\{\|V_{k}\|,\epsilon_{k}\}\leq 10^{-12}$, replacing the inexact criterion in (11); and an IPG method with a fixed Stepsize, corresponding to (Millan:2019, , Algorithm 2) with $\alpha_{k}=1/L$ and $\sigma=0.9$, where $L$ is the Lipschitz constant of $f$. This algorithm is denoted by IPG-FixStep and is defined as
$X_{k+1}=\tilde{X}_{k}-V_{k},\quad\forall k\geq 0,$
where $\tilde{X}_{k}$ and $V_{k}$ satisfy
$LV_{k}\in\partial_{\varepsilon_{k}}g(\tilde{X}_{k})+L(\tilde{X}_{k}-X_{k})+\nabla f(X_{k}),\quad\|V_{k}\|^{2}+\frac{2\varepsilon_{k}}{L}\leq 0.9\|X_{k}-\tilde{X}_{k}\|^{2}.$
The initialization parameters for the IPG-ELS method were set as $\tau=0.8$, $\theta=0.5$, $\gamma_{1}=\gamma_{2}=1.1$, and $\alpha=0.01$. For all tests, we initialized $X_{0}=0\in\mathbbm{R}^{n\times m}$, and set $\lambda_{\text{row}}=\lambda_{\text{col}}=0.01$. The IPG-ELS method was executed for $101$ outer iterations to establish a baseline performance metric, $F_{*}:=F(X_{101})$. The other two algorithms were terminated as soon as $F(x_{k})\leq F_{*}$. The algorithms were evaluated on six datasets from Cano:2005 ; Guyon:2004 ; Team:2008 : Colon tumor ($62\times 2000$), heart disease ($303\times 14$), central nervous system (CNS) ($60\times 7129$), lung cancer-Michigan ($96\times 7129$), Secom ($1567\times 590$), and Cina0 ($132\times 16033$).
Each matrix $W$ was normalized to have a unit Frobenius norm, with an additional step of centering each column. Subsequently, the resulting matrices were multiplied by a constant, which plays a crucial role in controlling the Lipschitz constant of the function $f$. The experiments were conducted using the Python programming language, which was installed on a machine equipped with a 3.5 GHz Dual-Core Intel Core i5 processor and 16 GB of 2400 MHz DDR4 memory.
In Tables 1 and 2, we report the Lipschitz constant of the gradient of $f$ (denoted as Lips), the number of outer iterations (O-IT), the number of inner iterations (I-IT), the number of linesearch iterations (LS-IT), and the total running time in seconds (Time). The results indicate that, in terms of CPU times, the IPG-ELS method outperforms the other two methods. This efficiency can be attributed to two main factors: (i) the PG-ELS method requires a significantly higher number of inner iterations to solve the proximal subproblem ”exactly”, and (ii) the IPG-FixStep method employs a small stepsize of $1/L$ in the gradient step.
Problem Lips Method $F(X_{k})$ O-IT I-IT LS-IT Time
Colon Tumor 41.58 IPG-ELS 1.1056 101 195 318 28.88
PG-ELS 1.1056 104 751 324 36.19
IPG-FixStep 1.1056 1450 1451 - 171.72
Colon Tumor 665.32 IPG-ELS 2.3647 101 101 823 32.74
PG-ELS 2.3623 128 455 1038 44.02
IPG-FixStep 2.3647 1103 1104 - 97.72
Colon Tumor 5133.69 IPG-ELS 5.8989 101 101 1134 110.86
PG-ELS 5.8832 119 328 1324 116.26
IPG-FixStep 5.8981 586 587 - 138.92
Heart Disease 77.12 IPG-ELS 0.1732 101 178 539 1.46
PG-ELS 0.1732 119 876 603 4.28
IPG-FixStep 0.1732 487 488 - 3.52
Heart Disease 1233.99 IPG-ELS 0.3129 101 101 903 0.85
PG-ELS 0.3119 89 368 788 1.71
IPG-FixStep 0.3520 2001 2001 - 11.85
Heart Disease 9521.59 IPG-ELS 0.4995 101 101 1222 0.99
PG-ELS 0.4992 227 622 2690 5.46
IPG-FixStep 0.7242 2001 2001 - 43.92
CNS 41.95 IPG-ELS 0.9519 101 182 341 364.44
PG-ELS 0.9518 119 714 396 501.59
IPG-FixStep 0.9519 1217 1218 - 1675.05
CNS 671.17 IPG-ELS 2.1153 101 101 768 544.91
PG-ELS 2.1119 148 725 1110 878.39
IPG-FixStep 2.2193 2001 2001 - 2749.17
CNS 5178.80 IPG-ELS 6.0665 101 101 1130 1089.70
PG-ELS 6.0613 102 343 1138 1169.25
IPG-FixStep 6.0655 775 776 - 1677.93
Table 1: Performance of the IPG-ELS, PG-ELS and IPG-FixStep methods on $3$ data sets.
Problem Lips Method $F(X_{k})$ O-IT I-IT LS-IT Time
Lung cancer 52.58 IPG-ELS 0.8985 101 179 443 422.07
PG-ELS 0.8984 105 837 457 521.46
IPG-FixStep 0.8985 678 679 - 962.80
Lung cancer 841.22 IPG-ELS 2.7632 101 101 845 591.92
PG-ELS 2.7623 131 768 1085 851.75
IPG-FixStep 2.7631 588 589 - 838.30
Lung cancer 2658.70 IPG-ELS 3.6391 101 101 992 696.61
PG-ELS 3.6378 161 740 1574 1173.05
IPG-FixStep 3.8819 2001 2001 - 2969.58
Secom 45.78 IPG-ELS 0.6438 101 175 373 66.64
PG-ELS 0.6438 99 9247 360 694.81
IPG-FixStep 0.6438 857 858 - 225.29
Secom 732.50 IPG-ELS 0.8587 101 101 779 82.81
PG-ELS 0.8586 102 4931 795 423.80
IPG-FixStep 0.8586 1662 1663 - 440.13
Secom 5652.06 IPG-ELS 1.6981 101 101 1138 104.54
PG-ELS 1.6899 125 1346 1388 215.08
IPG-FixStep 1.6977 801 802 - 211.93
Cina0 68.39 IPG-ELS 0.7972 101 245 487 3041.60
PG-ELS 0.7972 104 1490 483 5005.84
IPG-FixStep 0.7972 608 790 - 5530.44
Cina0 527.70 IPG-ELS 1.2817 101 250 838 3925.72
PG-ELS 1.2817 94 1789 693 5712.33
IPG-FixStep 1.2817 1010 1011 - 8741.06
Cina0 8443.20 IPG-ELS 3.6531 101 104 1126 3791.25
PG-ELS 3.6530 284 3083 3061 15338.18
IPG-FixStep 3.8493 2001 2001 - 17104.23
Table 2: Performance of the IPG-ELS, PG-ELS and IPG-FixStep methods on $3$ data sets.
6 Conclusions
In this work, we present an inexact proximal gradient method for solving composite convex optimization problems. This method features a novel explicit line search using the relative-type inexact solution of the proximal subproblem. Our approach is primarily designed to solve splitting problems when the objective function is the sum of differentiable and nondifferentiable convex functions, and the analytical computation of the proximal operator is not available. Notably, the convergence of the proposed method is established without assuming Lipschitz continuity of the gradient of the smooth function. This method addresses the need for a balance between computational efficiency and the accuracy of solving the proximal subproblem, a common challenge in practice.
We have confirmed the convergence and iteration complexity of our method, validating its theoretical soundness and practical utility. Numerical experiments demonstrate its applicability and efficiency. Our method maintains convergence rates while efficiently managing relative inexact solutions of the proximal operator. The numerical results indicate that the proposed method competes effectively with both the exact proximal gradient method and the inexact proximal gradient method with a fixed stepsize.
Acknowledgements.
YBC was partially supported by the NSF Grant DMS-2307328, and by an internal grant from NIU. MLNG was partially supported by CNPq Grants 405349/2021-1 and 304133/2021-3. JGM was partially supported by CNPq Grant 312223/2022-6.