### v2

Abstract
This paper presents and investigates an inexact proximal gradient method for solving composite convex optimization problems characterized by an objective function composed of a sum of a full-domain differentiable convex function and a non-differentiable convex function. We introduce an explicit line search applied specifically to the differentiable component of the objective function, requiring only a relative inexact solution of the proximal subproblem per iteration. We prove the convergence of the sequence generated by our scheme and establish its iteration complexity, considering both the functional values and a residual associated with first-order stationary solutions. Additionally, we provide numerical experiments to illustrate the practical efficacy of our method.
Keywords:
Linesearch Iteration complexity Nonsmooth and convex optimization problemProximal gradient method Relative error rule
MSC:
65K05 90C25 90C30
journal: JOTA
1 Introduction
In this paper, our focus is on addressing nonsmooth convex optimization problems characterized by the following formulation:
$\min\,F(x):=f(x)+g(x)\quad\text{subject to}\quad x\in\mathbbm{E},$ (1)
where $\mathbbm{E}$ represents a finite-dimensional Euclidean space. The function $g:\mathbbm{E}\to\overline{\mathbbm{R}}:=\mathbbm{R}\cup\{+\infty\}$ is a nonsmooth, proper, and lower semicontinuous convex function. The function $f:\mathbbm{E}\to\mathbbm{R}$ is a continuously differentiable and convex function. Throughout the paper, we denote the optimal value and solution set of problem (1) by $F_{*}$ and $S_{*}$, respectively. From now on, we assume that $S_{*}\neq\emptyset$.
Proximal gradient methods effectively solve optimization problems such as (1). The main step in the proximal gradient method involves evaluating the proximal operator ${\bf prox}_{g}:\mathbbm{E}\to{\bf dom\,}g:=\{x\in\mathbbm{E}\mid g(x)<+\infty\}$ defined as follows:
${\bf prox}_{g}(x):=\underset{y\in\mathbbm{E}}{\arg\min}\;\left\{g(y)+\frac{1}{2}\left\|y-x\right\|^{2}\right\},$ (2)
[…]
Proximal Gradient Method (PGM) Let $x_{0}\in{\bf dom\,}g$. Compute $\lambda_{k}>0$ and $\tilde{x}_{k}:={\bf prox}_{\lambda_{k}g}\left(x_{k}-\lambda_{k}\nabla f(x_{k})\right).$ (3) Choose $\beta_{k}\in(0,1]$ and compute $x_{k+1}:=x_{k}+\beta_{k}(\tilde{x}_{k}-x_{k}).$ (4)
The coefficients $\lambda_{k}$ and $\beta_{k}$, referred to as stepsizes, can be determined based on backtracking linesearch procedures. Such strategies are essential whenever the global $L$-Lipschitz continuity of the gradient of $f$ fails or even when computing an acceptable upper bound for $L$ is challenging. This situation is often encountered in numerous applications; for instance, in inverse problems based on non-Euclidean norms Bredies:2009 ; Schuster:2012 or Bregman distances such as the Kullback-Leibler divergence Salzo:2014 ; Bonettini:2012 ; Csiszar:1991 ; Vardi:1985 . Moreover, even when $L$ is known, linesearches may allow for longer steps toward the solution by using local information at every iteration.
There are several possible choices for these stepsizes, each impacting the algorithm’s performance in different ways; see, for instance, Salzo:2017a ; BelloCruz:2016a ; BelloCruz:2016c ; Bello-Cruz:2021 . It is important to note that in order to compute the stepsize $\lambda_{k}$ using a backtracking linesearch at each iteration $k$, the proximal operator may need to be evaluated multiple times within the procedure. Conversely, the stepsize $\beta_{k}$ can be selected by evaluating the proximal operator only once per iteration. In this context, we will refer to explicit linesearch to describe a backtracking procedure that determines $\beta_{k}$ after setting $\lambda_{k}$ as a constant for all $k$. This type of explicit strategy, first presented in BelloCruz:2016a , is particularly advantageous, especially in cases where evaluating the proximal operator is challenging. The function $g$ is often complex enough that the corresponding proximal operator lacks an analytical solution. In such cases, an ad-hoc algorithm should be employed to evaluate the proximal operator inexactly. For instance, consider $\mathbb{E}=\mathbbm{R}^{n}$, and let $g:\mathbbm{R}^{n}\to\mathbbm{R}$ be defined as
$g(x)=\|x\|_{1}+\lambda\sum_{i<j}\max\{|x_{i}|,|x_{j}|\},$
with $\lambda>0$, a form encountered in sparse regression problems with grouping, as discussed in Zhong:2012 . Similarly, in the context of matrix factorization, consider $\mathbb{E}=\mathbbm{R}^{n\times m}$ for the CUR-like factorization optimization problem Barre:2022 ; JMLR:v12:mairal11a , where the goal is to approximate a matrix $W\in\mathbbm{R}^{m\times n}$ with $X\in\mathbbm{R}^{n\times m}$ having sparse rows and columns. In this case, $g:\mathbbm{R}^{n\times m}\to\mathbbm{R}$ is given by
$g(X)=\lambda_{\text{row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+\lambda_{\text{col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2},$
where $\lambda_{\text{row}},\lambda_{\text{col}}>0$. This nondifferentiable term in problem (1) will be considered in the numerical illustrations of this paper. For further examples and discussions, see BelloCruz:2017a ; Bello-Cruz:2023 ; Salzo:2012 ; Villa:2013 ; Schmidt:2011 ; Jiang:2012 ; Millan:2019 ; Barre:2022 ; BelloCruz:2022b . Of course, there are some cases when the exact analytical solution of the proximal operator is available, such as when $g:\mathbbm{R}^{n}\to\mathbbm{R}$ as $g(x)=\|x\|_{1}$ or the indicator function of a simple convex and closed set, see, for instance, Beck:2009 ; Combettes:2005 .
Consequently, in practice, the evaluation of the proximal operator is often done inexactly. A basic inexactness criterion is to approximately evaluate the proximal operator using an exogenous sequence of error tolerances which, in general, must be summable in order to guarantee the convergence of the proximal iterations. Such a diminishing sequence is a priori chosen without using any information that may be available along with the iterations. Usually, the restrictive summability condition forces the solution of the proximal subproblem to be increasingly accurate, often more than necessary; see, for instance, Jiang:2012 ; Salzo:2012 ; Aujol:2015 ; Villa:2013 ; Schmidt:2011 . In the past two decades, relative error criteria have been considered an effective and practical way of controlling the inexactness in solving proximal subproblems of several algorithms, including FISTA, ADMM, augmented Lagrangian, Douglas-Rachford, and proximal gradient methods; refer to Eckstein:2013 ; Eckstein:2018 ; Alves:2020 ; Adona:2019 ; Adona:2020 ; Bonettini:2016 ; Bello-Cruz:2023 ; Millan:2019 for examples. Relative error criteria are often easy to verify in practice and have the advantage of exploiting the available information at a specific iteration, being, therefore, an interesting alternative to the aforementioned exogenous sequences.
In this paper, we propose and analyze an inexact proximal gradient method (PGM) for solving problem (1). We present a novel relative inexactness criterion for solving the proximal subproblems which somewhat resembles the ideas of relative error criteria introduced in Monteiro:2010 ; Solodov:1999a , but incorporates some new elements to control the objective function for the inexact solution. The proposed scheme requires only one inexact solution of the proximal subproblem per iteration, and the stepsizes are computed through a relaxed explicit linesearch procedure, applied specifically to $f$, that takes into account the residuals obtained from the proximal subproblem and enables the iteration to address non-Lipschitz optimization problems effectively. We show that the sequence generated by our method converges to a solution of problem (1). Moreover, we establish its iteration complexity in terms of both the function values and the residuals associated with an approximate stationary solution. We also present some numerical experiments to illustrate the performance of the proposed method.
It is worth mentioning that the PGM proposed in this paper can be regarded as an inexact version of the ones analyzed in Salzo:2017a ; BelloCruz:2016a , where explicit linesearches were studied in the PGM setting. On the other hand, an inexact version of the PGM with an explicit linesearch was also proposed in Bonettini:2016 . The authors developed a general framework based on variable metrics and established results on convergence rates for the sequence generated by their method. While, in the special case where the metric is fixed and chosen as the Euclidean one, their algorithm shares some similarities with ours, it employs different type of linesearch and inexact criteria for solving the proximal subproblem. Further details on the differences between our algorithm and the method in Bonettini:2016 are discussed in Subsection 3.3.
The paper is structured as follows: Section 2 presents definitions, basic facts, and auxiliary results. The concept of an approximate solution for the proximal subproblem, along with the description of algorithms that can be employed to compute it, is detailed in Section 3. In this latter section, we discuss some works that analyze relative inexact proximal solution criteria related to ours. Section 4 introduces the inexact proximal gradient method with an explicit linesearch (IPG-ELS) and establishes some of its fundamental properties. Section 5 analyzes the full convergence of the sequence generated by the IPG-ELS method and establishes its iteration complexity bounds in terms of functional values and a residual associated with the stationary condition for problem (1). Some numerical experiments illustrating the behavior of the proposed scheme are reported in Section 6. Finally, concluding remarks are provided in Section 7.
2 Preliminary
In this section, we present some preliminary material and notations that will be used throughout this paper.
Let $g:\mathbbm{E}\to\overline{\mathbbm{R}}$ be a proper, lower semicontinuous, and convex function. For a given $\varepsilon\geq 0$, the $\varepsilon$-subdifferential of $g$ at $x\in{\bf dom\,}g=\{x\in\mathbbm{E}\mid g(x)<+\infty\}$, denoted by $\partial_{\varepsilon}g(x)$, is defined as
$\partial_{\varepsilon}g(x):=\{v\in\mathbbm{E}\mid g(y)\geq g(x)+\langle v,y-x\rangle-\varepsilon,\ \forall y\in\mathbbm{E}\}.$ (5)
When $x\notin{\bf dom\,}g$, we define $\partial_{\varepsilon}g(x)=\emptyset$. Any element $v\in\partial_{\varepsilon}g(x)$ is called an $\varepsilon$-subgradient. If $\varepsilon=0$, then $\partial_{0}g(x)$ becomes the classical subdifferential of $g$ at $x$, denoted by $\partial g(x)$. It follows immediately from (5) that
$v\in\partial_{\varepsilon}g(y),\;u\in\partial g(x)\quad\text{implies}\quad\langle v-u,y-x\rangle\geq-\varepsilon.$ (6)
We present two useful auxiliary results for $\partial_{\varepsilon}g$. The first one is the closedness of the graph of $\partial_{\varepsilon}g$ and the second is the so-called transportation formula; see Propositions 4.1.1 and 4.2.2 of Hiriart-Urruty:1993 .
Proposition 1 (Closed Graph Property)
Let $(\varepsilon_{k},x_{k},v_{k})_{k\in\mathbbm{N}}\subseteq\mathbbm{R}_{+}\times\mathbbm{E}\times\mathbbm{E}$ be a sequence converging to $(\varepsilon,x,v)$ with $v_{k}\in\partial_{\varepsilon_{k}}g(x_{k})$ for all $k\in\mathbbm{N}$. Then, $v\in\partial_{\varepsilon}g(x)$.
Proposition 2 (Transportation Formula)
With $x$ and $y$ in ${\bf dom\,}g,$ let $v\in\partial g(y)$. Then $v\in\partial g_{\epsilon}(x)$ where $\epsilon=g(x)-g(y)-\langle{v},{x-y}\rangle\geq 0$.
We now introduce a concept of approximate stationary solution to problem (1), which can be seen as a specialization of the one presented in (Monteiro:2010, , Eq. 1). First, note that $\bar{x}$ is a solution to problem (1) if and only if $0\in\nabla f(\bar{x})+\partial g(\bar{x})$. The concept of approximate stationary solution relaxes this inclusion by introducing a residual $v$ and enlarging $\partial g$ using $\partial_{\varepsilon}g$.
Definition 1 ($\eta$-Approximate Stationary Solution)
Given $\eta>0$, an element $\tilde{x}\in{\bf dom\,}g$ is said to be an $\eta$-approximate stationary solution to problem (1) with a residual pair $(v,\varepsilon)$ if
$v\in\nabla f(\tilde{x})+\partial_{\varepsilon}g(\tilde{x}),\quad\max\{\|v\|,\varepsilon\}\leq\eta.$ (7)
Next, we recall the definition of quasi-Fejér convergence, which is an important and well-known tool for establishing full convergence of gradient and proximal point type methods; see, for instance, (Iusem-Quasi-Fejer, , Definition 1).
Definition 2 (Quasi-Fejér Convergence)
Let $S$ be a nonempty subset of $\mathbbm{E}$. A sequence $(x_{k})_{k\in\mathbbm{N}}\subseteq\mathbbm{E}$ is said to be quasi-Fejér convergent to $S$ if and only if, for every $x\in S$, there exists a summable sequence $(\delta_{k})_{k\in\mathbbm{N}}\subseteq\mathbbm{R}_{+}$ such that
$\|x_{k+1}-x\|^{2}\leq\|x_{k}-x\|^{2}+\delta_{k},\qquad\forall k\in\mathbbm{N}.$ (8)
The following result presents the main properties of quasi-Fejér convergent sequences; see (Bauschke:1996, , Theorem $2.6$).
Lemma 1 (Quasi-Féjer Convergence Properties)
If $(x_{k})_{k\in\mathbbm{N}}$ is quasi-Fejér convergent to $S$, then the following statements hold:
The sequence $(x_{k})_{k\in\mathbbm{N}}$ is bounded;
If an accumulation point $\bar{x}$ of $(x_{k})_{k\in\mathbbm{N}}$ belongs to $S$, then $(x_{k})_{k\in\mathbbm{N}}$ is convergent to $\bar{x}$.
We conclude the section with a basic inequality that will be used in the next sections.
Lemma 2
For any $x,y\in\mathbbm{E}$, we have $\|x+y\|^{2}\leq 2\|x\|^{2}+2\|y\|^{2}$.
3 Inexact proximal solutions
In this section, we introduce a concept of approximate solutions for the proximal gradient subproblem (3). We then describe how certain well-known algorithms compute these approximate solutions when the objective function exhibits a specific structure. Finally, we discuss the most closely related inexact proximal criteria proposed in the literature.
In the following, we introduce our concept of approximate proximal solution. First, recall that, given $x\in\mathbbm{E}$, the exact solution of subproblem (3) with $\lambda_{k}=1$ and $x_{k}=x$ consists of finding $\tilde{x}$ such that
$\tilde{x}={\rm arg}\!\min_{y\in\mathbbm{E}}\left\{\left\langle\nabla f(x),y-x\right\rangle+g(y)+\frac{1}{2}\|y-x\|^{2}\right\}.$ (9)
Equivalently, this corresponds to solving the following monotone inclusion problem:
$0\in\nabla f(x)+\partial g(\tilde{x})+\tilde{x}-x.$ (10)
The concept of approximate solution given below consists of relaxing the above inclusion by introducing a residual pair $(v,\varepsilon)$ that satisfies a specific mixed-relative error condition.
Definition 3
Let $\tau\in(0,1]$, $\gamma_{1}>1$, $\gamma_{2}\geq 1$, and $\alpha\in[0,1-\tau]$. Given a point $x\in\mathbb{E}$, we say that $\tilde{x}\in\mathbb{E}$ is an inexact proximal solution of (10) if there exists a residual pair $\RR(v,\varepsilon)\in\mathbb{E}\times\RR_{+}$ such that the triple $(\tilde{x},v,\varepsilon)$ satisfies the following conditions:
$v\in\nabla f(x)+\partial_{\varepsilon}g(\tilde{x})+\tilde{x}-x,$ (11)
$g(\tilde{x}-v)-g(\tilde{x})-\langle\nabla f(x),v\rangle$ $+\frac{(1+\gamma_{1})}{2}\|v\|^{2}+(1+\gamma_{2})\varepsilon\leq\frac{(1-\tau-\alpha)}{2}\|x-\tilde{x}\|^{2}.$ (12)
Let us now make some remarks about the above definition. First note that the inclusion in (11) relaxes (10) by introducing an approximate solution $\tilde{x}$ together with a residues pair $(v,\varepsilon)$. The inequality in (12) provides a mechanism for controlling the residual pair $(v,\varepsilon)$, the functional value of $g$ at the approximate solution $\tilde{x}$, and the angle between $\nabla f(x)$ and the residual $v$, all in terms of the distance between $\tilde{x}$ and the point $x$. Additionally, (11)–(12) ensure that both $\tilde{x}$ and $\tilde{x}-v$ lie within ${\bf dom\,}g$.
Now note that if $\tilde{x}$ is the exact solution of the monotone inclusion problem (10), we immediately have
$(\tilde{x},v,\varepsilon):=\big{(}{\bf prox}_{g}(x-\nabla f(x)),0,0\big{)}$
satisfies (11)-(12), for all $\tau\in(0,1]$, $\gamma_{1}>1$, $\gamma_{2}\geq 1$ and $\alpha\in[0,1-\tau]$. Note also that if $x$ is not an exact solution of (1), it cannot be the exact solution of (10) either. Hence, by considering a point $\tilde{x}$ that differs from $x$ but remains sufficiently close to the exact solution of (10), it is direct that the right-hand side of (12) is strictly positive because $01-\tau-\alpha>$ 0. Consequently, this inequality will eventually be satisfied by a residual pair $(v,\varepsilon)$ that approaches zero, since its left-hand side also converges to zero in this case.
It is worth pointing out that depending on the structure of the problem and/or the algorithm utilized for computing an inexact solution of the proximal subproblem, the residual $v$ as in Definition 3 can be set as zero. In this case, (11)-(12) are equivalent to
$0\in\nabla f(x)+\partial_{\varepsilon}g(\tilde{x})+\tilde{x}-x,$ (13)
$\varepsilon\leq\frac{(1-\tau-\alpha)}{2(1+\gamma_{2})}\|x-\tilde{x}\|^{2}.$ (14)
In the following two subsections, we focus on a specific structure of the nonsmooth component function $g$ and provide a detailed discussion on how to compute a triple $(\tilde{x},v,\varepsilon)$ as described in Definition 3. Specifically, we address two scenarios: (i) $g$ has a separable structure, which includes applications such as CUR-like factorization problems; and (ii) $g$ is the indicator function of a special nonempty convex and compact set $C$. We also note that inexact proximal solutions for the $H$-weighted nearest correlation matrix problem and the convex regularized problem can be obtained from (Bello-Cruz:2023, , Section 3) through straightforward adaptations of the IR rules discussed therein.
3.1 Inexact proximal solutions for separable functions
In this subsection, we discuss how inexact proximal solutions of (10) as in Definition 3 can be computed if the nonsmooth component function $g$ has a separable structure of the form $g=g_{1}+g_{2},$ where both ${\bf prox}_{g_{1}}$ and ${\bf prox}_{g_{2}}$ admit closed-form solutions. In this case, the proximal subproblem corresponds to
$\min_{y\in\mathbbm{E}}\left\{g_{1}(y)+g_{2}(y)+\frac{1}{2}\|y-z\|^{2}\right\},$ (15)
where $z=x-\nabla f(x)$.
We next discuss how an inexact proximal solution to (15) can be obtained using either the Dykstra-like algorithm (Bauschke:2008c, , Theorem 3.3) or the alternating direction method of multipliers (ADMM) Alves:2020 ; Adona:2019 ; Adona:2020 . The approach related to the Dykstra algorithm is inspired by Schmidt:2011 (see also Barre:2022 ; Zhou:2022 ).
Let us first discuss the Dykstra-like algorithm. This algorithm, applied to problem (15) with initial points $z_{0}=z$, $p_{0}=0$, and $q_{0}=0$, generates the sequences:
$y_{\ell}$ $={{\bf prox}}_{g_{1}}(z_{\ell}+p_{\ell}),\quad$ $p_{\ell+1}$ $=z_{\ell}+p_{\ell}-y_{\ell},$ (16)
$z_{\ell+1}$ $={{\bf prox}}_{g_{2}}(y_{\ell}+q_{\ell}),\quad$ $q_{\ell+1}$ $=y_{\ell}+q_{\ell}-z_{\ell+1}.$
From the definition of $y_{\ell}$ in (16), it follows that $z_{\ell}+p_{\ell}-y_{\ell}\in\partial g_{1}(y_{\ell})$. This observation, combined with Proposition 2, yields
$z_{\ell}+p_{\ell}-y_{\ell}\in\partial_{\epsilon_{\ell}}g_{1}(z_{\ell+1}),\quad\mbox{where}\quad\epsilon_{\ell}=g_{1}(z_{\ell+1})-g_{1}(y_{\ell})-\langle{z_{\ell}+p_{\ell}-y_{\ell}},{z_{\ell+1}-y_{\ell}}\rangle\geq 0.$ (17)
On the other hand, it follows from the definition of $z_{\ell+1}$ in (16) that $y_{\ell}+q_{\ell}-z_{\ell+1}\in\partial g_{2}(z_{\ell+1}).$ Thus, by adding the inclusion in (17) and the latter one, we have
$z_{\ell}+p_{\ell}+q_{\ell}-z_{\ell+1}\in\partial_{\epsilon_{\ell}}g(z_{\ell+1}).$
Rearranging the terms and using that $z=x-\nabla f(x)$, we get
$z_{\ell}+p_{\ell}+q_{\ell}-z\in\partial_{\epsilon_{\ell}}g(z_{\ell+1})+z_{\ell+1}-x+\nabla f(x).$
It follows from the definitions of $p_{\ell+1}$ and $q_{\ell+1}$ in (16) that $p_{\ell+1}+q_{\ell+1}+z_{\ell+1}=p_{\ell}+q_{\ell}+z_{\ell}$ for all $\ell$. Then, from the latter inclusion we have
$z_{0}+p_{0}+q_{0}-z\in\partial_{\epsilon_{\ell}}g(z_{\ell+1})+z_{\ell+1}-x+\nabla f(x).$
Since $z=p_{0}+q_{0}+z_{0}$, the inclusion becomes
$0\in\partial_{\epsilon_{\ell}}g(z_{\ell+1})+z_{\ell+1}-x+\nabla f(x).$
Finally, if the Dykstra-like algorithm is terminated when
$\varepsilon_{\ell}\leq\frac{(1-\tau-\alpha)}{2(1+\gamma_{2})}\|x-z_{\ell}\|^{2},$ (18)
then $(\tilde{x},v,\varepsilon):=(z_{\ell},0,\varepsilon_{\ell})$ satisfies the conditions in (11)–(12).
We also note that, as shown in (Bauschke:2008c, , Theorem 3.3), the sequences $(y_{\ell})_{\ell\in\mathbbm{N}}$ and $(z_{\ell})_{\ell\in\mathbbm{N}}$ converge to the exact solution $\bar{x}$ of (15). Consequently, the sequence $(\varepsilon_{\ell})_{\ell\in\mathbbm{N}}$ converges to zero provided that the subdifferential of $g_{1}$ is bounded over a compact set, which holds for instance if ${\bf dom\,}g_{1}$ is an open set. Furthermore, as long as $\|x-\bar{x}\|>0$, it is guaranteed that the Dykstra-like algorithm will find a triple $(\tilde{x},0,\varepsilon)$ that satisfies the conditions in (11)–(12), since the left-hand side of (18) will converge to zero while the right-hand side remains strictly positive if $01-\tau-\alpha>$ 0.
Let us now discuss the ADMM. Note that (15) can be rewritten as
$\min\left\{\hat{g}_{1}(x)+\hat{g}_{2}(y):x-y=0\right\},$ (19)
where $\hat{g}_{1}(x)=g_{1}(x)$ and $\hat{g}_{2}(y)=g_{2}(y)+\frac{1}{2}\|y-z\|^{2}$. By applying the standard ADMM to the above problem, with initial point $(y_{0},\lambda_{0})\in\mathbbm{E}\times\mathbbm{E}$ and penalty parameter $c>0$, the following sequences are generated:
$x_{\ell}={\bf prox}_{g_{1}/c}\left(y_{\ell-1}+\frac{\lambda_{\ell-1}}{c}\right),\quad y_{\ell}={\bf prox}_{g_{2}/(c+1)}\left(\frac{z+cx_{\ell}-\lambda_{\ell-1}}{c+1}\right),\quad\lambda_{\ell}=\lambda_{\ell-1}-c(x_{\ell}-y_{\ell}).$ (20)
From the definition of $x_{\ell}$ in (20), it follows that $cy_{\ell-1}+\lambda_{\ell-1}-cx_{\ell}\in\partial g_{1}(x_{\ell})$. Combining the last inclusion with Proposition 2, we obtain
$cy_{\ell-1}+\lambda_{\ell-1}-cx_{\ell}\in\partial_{\varepsilon_{\ell}}g_{1}(y_{\ell}),\quad\text{where}\quad\varepsilon_{\ell}=g_{1}(y_{\ell})-g_{1}(x_{\ell})-\langle cy_{\ell-1}+\lambda_{\ell-1}-cx_{\ell},y_{\ell}-x_{\ell}\rangle\geq 0.$
Similarly, from the definition of $y_{\ell}$ in (20), we have $z+cx_{\ell}-\lambda_{\ell-1}-(c+1)y_{\ell}\in\partial g_{2}(y_{\ell})$. Consequently, from the above inclusions, it follows that
$c(y_{\ell-1}-y_{\ell})-y_{\ell}+z\in\partial_{\varepsilon_{\ell}}g(y_{\ell}),$
which, combined with the definition of $z$, yields
$v_{\ell}:=c(y_{\ell-1}-y_{\ell})\in\partial_{\varepsilon_{\ell}}g(y_{\ell})+y_{\ell}-x+\nabla f(x).$
Finally, if the ADMM is terminated when
$g(y_{\ell}-v_{\ell})-g(y_{\ell})-\langle\nabla f(x),v_{\ell}\rangle+\frac{(1+\gamma_{1})}{2}\|v_{\ell}\|^{2}+(1+\gamma_{2})\varepsilon_{\ell}\leq\frac{(1-\tau-\alpha)}{2}\|x-y_{\ell}\|^{2},$ (21)
then $(\tilde{x},v,\varepsilon):=(y_{\ell},v_{\ell},\varepsilon_{\ell})$ satisfies the conditions in (11)–(12).
Note that, since the sequence $((x_{\ell},y_{\ell},\lambda_{\ell}))_{\ell\in\mathbbm{N}}$ converges to $(\bar{x},\bar{y},\bar{\lambda})$ which satisfies $0\in\partial g_{1}(\bar{x})-\bar{\lambda}$, $0\in\partial g_{2}(\bar{y})+\bar{y}-z-\bar{\lambda}$ and $\bar{x}=\bar{y}$, it follows that $(v_{\ell})_{\ell\in\mathbbm{N}}$ converges to zero, and $(\varepsilon_{\ell})_{\ell\in\mathbbm{N}}$ also converges to zero, as long as the subdifferential of $g_{1}$ is bounded over a compact set. Furthermore, as long as $\|x-\bar{x}\|>0$, the ADMM is guaranteed to reach a triple $(\tilde{x},v,\varepsilon)$ satisfying the conditions in (11)–(12), since the left-hand side of (21) will converge to zero while the right-hand side remains strictly positive provided that $01-\tau-\alpha>$ 0.
3.2 Inexact projection computed by the Frank-Wolfe method
In this subsection, we assume that the nonsmooth component function $g$ is the indicator function of a nonempty, convex and compact set $C$. In this context, the proximal subproblem (9) (or equivalently the proximal monotone inclusion (10)) corresponds to project the point $x-\nabla f(x)$ onto the set $C$, and hence, the inexact proximal solution as in Definition 3 can be seen as an inexact projection of $x-\nabla f(x)$ onto the set $C$. By considering the case in which the residual $v$ can be set as $v=0$ (see (13)-(14)), we have that $\tilde{x}\in C$ is an inexact projection with residual $\RR(0,\varepsilon)\in\mathbb{E}\times\RR_{+}$ if and only if
$\langle x-\nabla f(x)-\tilde{x},y-\tilde{x}\rangle\leq\varepsilon\leq\frac{(1-\tau-\alpha)}{2(1+\gamma_{2})}\|x-\tilde{x}\|^{2},\quad\forall y\in C.$ (22)
For example, one can use the conditional gradient (CondG) method, a.k.a. Frank-Wolfe method Frank:1956 ; Jaggi:2013 , to compute $\tilde{x}\in C$ as in (22). Indeed, given $z_{\ell}\in C$, the $\ell$-th step of the CondG method, applied to solve the projection problem
$\min_{y\in C}\frac{1}{2}\|y-x+\nabla f(x)\|^{2},$ (23)
first finds $w_{\ell}$ as the minimizer of the linear function $\langle{z_{\ell}-x+\nabla f(x)},{\cdot-z_{\ell}}\rangle$ over $C$ and then sets $z_{\ell+1}=(1-\alpha_{\ell})z_{\ell}+\alpha_{\ell}w_{\ell}$, where $\alpha_{\ell}:=\min\{1,\langle{x-\nabla f(x)-z_{\ell}},{w_{\ell}-z_{\ell}}\rangle/\|z_{\ell}-w_{\ell}\|^{2}\}$. If, at iteration $\ell$, the point $z_{\ell}$ and $w_{\ell}$ satisfy the stopping criterion:
$\varepsilon_{\ell}:=\langle x-\nabla f(x)-z_{\ell},w_{\ell}-z_{\ell}\rangle\leq\frac{(1-\tau-\alpha)}{2(1+\gamma_{2})}\|x-z_{\ell}\|^{2},$ (24)
then $\tilde{x}:=z_{\ell}$ is an inexact projection onto $C$ with residual $(v,\varepsilon):=(0,\varepsilon_{\ell}).$ Furthermore, since the objective function in (23) is strongly convex, the sequences $(z_{\ell})_{\ell\in\mathbbm{N}}$ and $(w_{\ell})_{\ell\in\mathbbm{N}}$ converge to $P_{C}(x-\nabla f(x))$. Hence, as long as $\|x-P_{C}(x-\nabla f(x))\|>0$ and $01-\tau-\alpha>$ 0, the conditional gradient method is guaranteed to reach a triple $(\tilde{x},0,\varepsilon)$ satisfying the conditions in (13)–(14). This is ensured because the left-hand side of (24) converges to zero, while the right-hand side remains strictly positive.
3.3 Related concepts of inexact proximal solutions
In this subsection, we discuss some works that analyze relative inexact proximal solution criteria, which are related to the concept introduced in Definition 3. Specifically, we review the criteria proposed in Solodov:1999a ; Millan:2019 ; Bonettini:2016 .
We start by recalling the error criterion introduced in Solodov:1999a , as applied to the proximal subproblem (2). It is worth noting that this criterion was originally proposed in the broader context of monotone inclusions. For a given point $x\in\mathbb{E}$, the proximal subproblem with a prox-stepsize $\lambda>0$ consists of computing $\tilde{x}$ such that
$0\in\lambda\partial g(\tilde{x})+\tilde{x}-x.$
The criterion in Solodov:1999a relaxes this inclusion by finding a triple $\RR(\tilde{x},w,\varepsilon)\in\mathbb{E}\times\mathbb{E}\times\RR_{+}$ that satisfies
$w\in\partial_{\varepsilon}g(\tilde{x}),\qquad\|\lambda w+\tilde{x}-x\|^{2}+2\lambda\varepsilon\leq\sigma\|\tilde{x}-x\|^{2},$ (25)
for some scalar $\sigma\in[0,1)$.
The hybrid proximal extragradient (HPE) method proposed in Solodov:1999a consists mainly of two basic steps: first, it computes an approximate proximal solution based on (25), and then an extragradient step is performed to determine the next iterate, given by $x^{+}=x-\lambda w$. The convergence properties of the HPE method were analyzed in Solodov:1999a , while its iteration complexity was established in Monteiro:2010 . As demonstrated in Solodov:1999a ; Monteiro:2010 , the HPE method can be seen as a general framework that encompasses various well-known algorithms, including the Korpelevich method Korpelevich:1976 and Tseng’s modified forward-backward splitting (Tseng-MFBS) method Tseng:2000 , among others.
Tseng-MFBS method can be, in particular, applied for solving (1). In this case, it generates a sequence $(x_{k})_{k\in\mathbbm{N}}$ as follows:
$x_{k+1}=\bar{x}_{k}-\lambda(\nabla f(\bar{x}_{k})-\nabla f(x_{k})),\qquad\bar{x}_{k}:={\bf prox}_{\lambda g}(x_{k}-\lambda\nabla f(x_{k})),$ (26)
where $x_{0}$ is a given initial point, $\lambda>0$ is the stepsize, and ${\bf prox}_{\lambda g}$ is defined in (2).
As shown in Monteiro:2010 ; Solodov:1999a , this scheme is an instance of the HPE method when $\nabla f$ is $L$-Lipschitz continuous. Indeed, by defining
$w_{k}:=\frac{1}{\lambda}(x_{k}-\bar{x}_{k})+\nabla f(\bar{x}_{k})-\nabla f(x_{k}),$
it can be verified that $(\tilde{x},w,\varepsilon):=(\bar{x}_{k},w_{k},0)$ satisfies the conditions in (25) with $\sigma=\lambda/L$.
Most recently, Millan:2019 proposed inexact PGMs for solving (1), where both component functions $f$ and $g$ may be nonsmooth. Algorithm 2 in this reference (with a fixed stepsize $\alpha_{k}=\lambda$ and $C=\mathbb{E}$) applied to (1), with the additional restriction that ${\bf dom\,}g=\mathbb{E}$, generates a sequence $(x_{k})_{k\in\mathbbm{N}}$ according to the following scheme: given $x_{k}\in\mathbb{E}$ and $\sigma\in[0,1)$, first let $z_{k}:=x_{k}-\lambda\nabla f(x_{k})$, and then compute a triple $\RR(\tilde{x}_{k},w_{k},\varepsilon_{k})\in\mathbb{E}\times\mathbb{E}\times\RR_{+}$ such that
$w_{k}\in\partial_{\varepsilon_{k}}g(\tilde{x}_{k}),\qquad\|\lambda w_{k}+\tilde{x}_{k}-z_{k}\|^{2}+2\lambda\varepsilon_{k}\leq\sigma\|\tilde{x}_{k}-z_{k}\|^{2},$
and update $x_{k}$ as $x_{k+1}=z_{k}-\lambda w_{k}$. Note that, by defining $v_{k}=\lambda w_{k}+\tilde{x}_{k}-z_{k}$, the above conditions, combined with the definition of $z_{k}$, are equivalent to
$\frac{1}{\lambda}v_{k}\in\nabla f(x_{k})+\partial_{\varepsilon_{k}}g(\tilde{x}_{k})+\frac{1}{\lambda}(\tilde{x}_{k}-x_{k}),\qquad\|v_{k}\|^{2}+2\lambda\varepsilon_{k}\leq\sigma^{2}\|\tilde{x}_{k}-z_{k}\|^{2}.$
Note that, the above inclusion corresponds to the one in (11) if $\lambda=1$. On the other hand, while the latter inequality represents a relative inexact proximal criterion, it differs from the one in (12). It is worth noting that Millan:2019 did not incorporate any linesearch procedure. Instead, they considered stepsizes such as fixed, exogenous, and Polyak-type.
We end this subsection by discussing Bonettini:2016 , which also studies a PGM with inexact proximal solutions and an explicit linesearch. To simplify the discussion, we assume that the prox-parameter in Bonettini:2016 is constant and equal to one, and that the variable metric, used to define their proximal subproblem, is fixed and coincides with the Euclidean one. In this case, at the $k$-th iteration, they compute a pair $\RR(\tilde{x}_{k},\varepsilon_{k})\in\mathbb{E}\times\RR_{+}$ such that
$0\in\nabla f(x_{k})+\partial_{\varepsilon_{k}}g(\tilde{x}_{k})+\tilde{x}_{k}-x_{k},$ (27)
i.e., the triple $(\tilde{x}_{k},0,\varepsilon_{k})$ is a solution of (11) (with a null residual $v_{k}$). Two types of conditions are assumed to control the residual sequence $(\varepsilon_{k})_{k\in\mathbbm{N}}$: (i) the summable absolute error condition; and (ii) a relative error condition. We focus our attention on the second one, which is more closely related to our inexact criterion. The authors assume that the pair $(\tilde{x}_{k},\varepsilon_{k})$ satisfies (27) and the following conditions:
$\varepsilon_{k}\leq-t\tilde{h}_{s}(\tilde{x}_{k}),\qquad\tilde{h}_{s}(\tilde{x}_{k})<0,$ (28)
for some $s\in(0,1]$ and $t>0$, where
$\tilde{h}_{s}(u)=\langle\nabla f(x_{k}),u-x_{k}\rangle+g(u)-g(x_{k})+\frac{s}{2}\|u-x_{k}\|^{2}.$ (29)
Note that $\tilde{h}_{s}$ is closely related to the objective function of the subproblem (9) and becomes negative at its unique solution. In particular, the second condition in (28) holds for some $\tilde{x}_{k}$ sufficiently close to the exact solution of (9).
We claim that if $(\tilde{x},v,\varepsilon):=(\tilde{x}_{k},0,\varepsilon_{k})$ is an approximate proximal solution in the sense of Definition 3, then the pair $(\tilde{x}_{k},\varepsilon_{k})$ satisfies (27)-(28) for some $t>0$. Indeed, in this case, (11) becomes
$0\in\nabla f(x)+\partial_{\varepsilon}g(\tilde{x})+\tilde{x}-x$
and (12) is equivalent to
$\varepsilon\leq\frac{1-\tau-\alpha}{2(1+\gamma_{2})}\|x-\tilde{x}\|^{2}.$ (30)
The above inclusion, together with the definitions of $\tilde{h}_{s}$ in (29) and the $\varepsilon$-subdifferential in (5), imply that
$\tilde{h}_{s}(\tilde{x})+\left(\frac{2-s}{2}\right)\|x-\tilde{x}\|^{2}\leq\varepsilon.$
Combining the latter two inequalities, we obtain
$\tilde{h}_{s}(\tilde{x})\leq\left[\frac{1-\tau-\alpha}{2(1+\gamma_{2})}-\frac{2-s}{2}\right]\|x-\tilde{x}\|^{2}.$
Hence, if $t$ is a scalar such that
$t\geq\frac{1-\tau-\alpha}{(2-s)(1+\gamma_{2})+\tau+\alpha-1}.$ (31)
Thus, in view of the latter inequality and (30) that
$\varepsilon\leq\frac{1-\tau-\alpha}{2(1+\gamma_{2})}\|x-\tilde{x}\|^{2}\leq-t\tilde{h}_{s}(\tilde{x}),$
which implies that the pair $(\tilde{x},\varepsilon)$ satisfies the conditions in (28) with $t$ as in (31), proving the claim.
We conclude by observing that even if the residue $v$ in Definition 3 is chosen as null, then the IPG-ELS algorithm differs from the scheme proposed inBonettini:2016 . This is because the linesearch we propose is applied only to the smooth part $f$ of the objective function $F$, whereas the Armijo-type linesearch procedure considered in Bonettini:2016 is applied to the whole objective function $F$. Moreover, if $v\neq 0$, then our inexact criterion differs significantly from the one in Bonettini:2016 . In addition, our linesearch procedure accounts for not only the proximal inexact solution $\tilde{x}$ but also the residue pair $(v,\varepsilon)$. It is also worth mentioning that the latter reference did not analyze the iteration complexity to achieve an approximate stationary solution in the sense of Definition 1. Finally, Bonettini:2016 also proposes a second criterion, where an approximate proximal solution $\tilde{x}$ of the subproblem (10) is accepted if
$\tilde{h}_{1}(\tilde{x})\leq\eta\tilde{h}_{1}(\bar{x}),$
where $\eta\in(0,1]$ and $\bar{x}$ being is exact solution of (10).
4 Inexact Proximal Gradient Method
In this section, we introduce the inexact proximal gradient method with an explicit linesearch and establish some basic properties.
In the following, we formally describe the proposed method.
[…]
Remark 1 (The Explicit Lineasearch)
Note that the novel explicit linesearch in Step 4 of the IPG-ELS is related to the ones proposed in Salzo:2017a ; BelloCruz:2016a when $v_{k}=0$ and $\varepsilon_{k}=0$. Moreover, note that the linesearch does not evaluate the inexact proximal subproblem inside the inner loop of Step 4. Hence, only one inexact proximal solution is computed per iteration.
We introduce a useful lemma that plays a crucial role in analyzing the stopping criterion of the IPG-ELS method.
Lemma 3 (Iteration Inequality Condition)
The following inequality holds for every iteration $k$.
$\left\langle\tilde{x}_{k}-x_{k},v_{k}\right\rangle+\frac{(\gamma_{1}-1)}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}\leq\frac{(1-\tau-\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2}.$ (35)
Proof It follows from (32) that
$v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k})\in\partial_{\varepsilon_{k}}g(\tilde{x}_{k}).$
Now using (5) together with the fact that $y_{k}=\tilde{x}_{k}-v_{k}$, we have
$g(y_{k})-g(\tilde{x}_{k})$ $\geq\left\langle v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k}),y_{k}-\tilde{x}_{k}\right\rangle-\varepsilon_{k}$
$=\left\langle v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k}),-v_{k}\right\rangle-\varepsilon_{k},$
which is equivalent to
$\Gamma_{k}:=g(y_{k})-g(\tilde{x}_{k})-\langle\nabla f({x}_{k}),v_{k}\rangle$ $\geq-\|v_{k}\|^{2}+\langle\tilde{x}_{k}-x_{k},v_{k}\rangle-\varepsilon_{k}.$ (36)
On the other hand, considering the definitions of $\Gamma_{k}$ and $y_{k}$, we observe that (33) is equivalent to
$\Gamma_{k}+\frac{(1+\gamma_{1})}{2}\|v_{k}\|^{2}+(1+\gamma_{2})\varepsilon_{k}\leq\frac{(1-\tau-\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2},$
which, combined with (36), yields the desired inequality (35). ∎
The following result demonstrates that the termination criterion of the IPG-ELS method, as specified in Step 3, is satisfied only when a solution to problem (1) is identified.
Lemma 4 (Stopping at a Solution)
The IPG-ELS method terminates at the $k$-th iteration if and only if $x_{k}$ is a solution to problem (1).
Proof Assume that the IPG-ELS method stops at the $k$-th iteration. In view of Step 3, we have $\tilde{x}_{k}=x_{k}$. Hence, it follows from (35) that
$\frac{(\gamma_{1}-1)}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}\leq 0.$
Since $\gamma_{1}>1$ and $\gamma_{2}\geq 1$, we obtain $v_{k}=0$ and $\varepsilon_{k}=0$. Hence, in view of (32), we get that $0\in\nabla f(x_{k})+\partial g(x_{k}),$ concluding that $x_{k}$ is a solution of problem (1).
Assume now that $x_{k}$ is a solution of problem (1). Thus, $-\nabla f(x_{k})\in\partial g(x_{k}).$ It follows from (32) that $v_{k}+x_{k}-\tilde{x}_{k}-\nabla f(x_{k})\in\partial_{\varepsilon_{k}}g(\tilde{x}_{k})$. So, the $\varepsilon$-monotonicity of $\partial_{\varepsilon}g$ in (6) implies
$\langle v_{k}+x_{k}-\tilde{x}_{k},\tilde{x}_{k}-x_{k}\rangle\geq-\varepsilon_{k},$
which is equivalent to
$\langle v_{k},\tilde{x}_{k}-x_{k}\rangle\geq\|x_{k}-\tilde{x}_{k}\|^{2}-\varepsilon_{k}.$ (37)
Hence, it follows from (35) that
$\frac{(1+\tau+\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{(\gamma_{1}-1)}{2}\|v_{k}\|^{2}+(\gamma_{2}-1)\varepsilon_{k}\leq 0.$
Since $\gamma_{1}>1$, $\gamma_{2}\geq 1$, $\alpha\geq 0$, and $\tau>0$, we conclude that $x_{k}=\tilde{x}_{k}$ and $v_{k}=0$. Hence, from (33), we also have $\varepsilon_{k}=0$. Therefore, the IPG-ELS method stops at the $k$-th iteration. ∎
In the following, we establish the well-definedness of the linesearch procedure in Step 4 of the IPG-ELS method.
Lemma 5 (Finite Linesearch Termination)
The linesearch procedure in Step 4 stops after a finite number of steps.
Proof In view of Step 3, we have $x_{k}\neq\tilde{x}_{k}$. Assume, for the sake of contradiction, that the linesearch procedure does not terminate after a finite number of steps. Thus, for all $\beta\in\{1,\theta,\theta^{2},\ldots\}$, we have
$f\left(x_{k}+\beta(y_{k}-x_{k})\right)>f(x_{k})+\beta\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{\beta\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\beta\|v_{k}\|^{2}+\beta\gamma_{2}\varepsilon_{k},$
or, equivalently,
$\frac{f\left(x_{k}+\beta(y_{k}-x_{k})\right)-f(x_{k})}{\beta}-\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle>\frac{\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}.$
Given that $f$ is differentiable, the left-hand side of the above inequality approaches zero as $\beta\downarrow 0$, leading to the conclusion that
$0\geq\frac{\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}.$
This implies $x_{k}=\tilde{x}_{k}$ contradicting the assumption that $x_{k}\neq\tilde{x}_{k}$. ∎
The above lemma ensures that the linesearch procedure in Step 4 of the IPG-ELS method terminates after a finite number of iterations. This result is fundamental to the convergence analysis of the method, as it guarantees that the linesearch procedure is well-defined.
Next, we provide a remark on the possibility of initiating the linesearch with a larger stepsize.
Remark 2 (Over-relaxation Strategy)
In the current implementation, the linesearch algorithm begins with an initial stepsize $\beta=1$. Considering larger initial values for $\beta$, thereby adopting an over-relaxation strategy, may potentially accelerate convergence by extrapolating beyond standard update steps. This approach is related to the concept of conic averagedness BartzDaoPhan , which offers a framework for analyzing the convergence properties of fixed-point algorithms under relaxed conditions. However, initiating the linesearch with $\beta>1$ presents significant challenges to our current approach. While the linesearch remains well-defined for any $\beta>1$, and the proof of finite termination would be directly extended, the convergence and complexity analysis of our scheme depend on the convexity of the function $f$ and the appropriateness of the direction $y_{k}-x_{k}$ to ensure a sufficient decrease in the objective function. Therefore, our convergence analysis does not cover such cases, and the practical implications for this approach are unclear.
The subsequent analysis investigates the complexity of the linesearch procedure introduced in Step 4 of the IPG-ELS method. Here we assume that the gradient of the function $f$, $\nabla f$, is Lipschitz continuous only to establish an upper bound for the number of iterations required by the linesearch procedure in Step 4.
Lemma 6 (Lipschitz Condition and Linesearch Complexity)
Assume that $f$ has an $L$-Lipschitz continuous gradient and that $x_{k}$ is not a solution to problem (1). Then, any $\beta\leq\tau/(2L)$ satisfies (34). As a consequence, the linesearch procedure in Step 4 of the IPG-ELS method stops in at most
$\ell:=\left\lceil\frac{\ln(\min\{\tau/(2L),1\})}{\ln(\theta)}\right\rceil$ (38)
iterations.
Proof Since $\nabla f$ is $L$-Lipschitz continuous, for any $\beta>0$, we have
$f(x_{k}+\beta(y_{k}-x_{k}))-f(x_{k})-\beta\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle$ $\leq\frac{L\beta^{2}}{2}\|y_{k}-x_{k}\|^{2}.$
Hence, if $\beta\leq\tau/(2L)$, we conclude that
$f(x_{k}+\beta(y_{k}-x_{k}))-f(x_{k})-\beta\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle$ $\leq\frac{\tau\beta}{4}\|y_{k}-x_{k}\|^{2}=\frac{\tau\beta}{4}\|x_{k}-\tilde{x}_{k}+v_{k}\|^{2}$
$\leq\frac{\beta\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\tau}{2}\beta\|v_{k}\|^{2},$
using Lemma 2 in the last inequality. Since $\gamma_{1}>1>\tau$ and $\gamma_{2}\geq 0$, we have that (34) holds, thereby proving the first statement of the lemma. The last statement follows from the first one, given that the natural number $\ell$, defined in (39), satisfies $\beta_{\ell}:=\theta^{\ell}\leq\min\{\tau/(2L),1\}$. ∎
This lemma provides a sufficient condition to ensure that the lower bound of the sequence generated by our linesearch is strictly greater than 00. Specifically, if $\nabla f$ is $L$-Lipschitz continuous, then the stepsizes $\beta_{k}$, produced through the linesearch (34), are guaranteed to be bounded below by a positive constant $\beta>0$, i.e., $\beta_{k}\geq\beta$ for all $k\in\mathbb{N}$. Moreover, it is possible to relax the global Lipschitz condition to something local, such as $\nabla f$ being locally Lipschitz continuous around any solution of problem (1), as was done in Proposition 5.4(ii) of BelloCruz:2016a . In fact, the proof of Lemma 6 may be readily adapted to establish the same complexity now with respect to the locally Lipschitz constant $\mathcal{L}$.
Lemma 7 (Locally Lipschitz Condition and Linesearch Complexity)
If $\nabla f$ is $\mathcal{L}$-locally Lipschitz continuous at any solution of problem (1), then any $\beta\leq\tau/(2\mathcal{L})$ satisfies (34). Consequently, the linesearch procedure in Step 4 of the IPG-ELS method terminates after at most
$\ell:=\left\lceil\frac{\ln(\min\{\tau/(2\mathcal{L}),1\})}{\ln(\theta)}\right\rceil$ (39)
iterations.
Note that the assumption of the gradient of $f$ in problem (1) being locally, rather than globally, Lipschitz continuous is commonly encountered in practice. For instance, this condition arises in the Poisson linear inverse regularization problem with Kullback–Leibler divergence Csiszar:1991 ; Vardi:1985 . We emphasize that this weaker assumption is required exclusively for analyzing the complexity of the linesearch procedure, as it ensures that the stepsizes generated by the proposed linesearch are uniformly bounded away from zero. Additionally, it may be used as sufficient condition for establishing the convergence rate of the functional value sequence. It is worth noting that the finite termination of the linesearch and convergence of the IPG-ELS method do not rely on such assumption.
5 Convergence and Complexity Analysis of the IPG-ELS Method
In this section, we focus on analyzing the convergence properties of the IPG-ELS method. We establish the convergence and its iteration complexity in terms of functional values and a residual associated with an approximate solution, as defined in Definition 1.
We begin this section by presenting a result that is fundamental for establishing the convergence and the iteration complexity of the IPG-ELS method.
Proposition 3 (Key Inequality for the IPG-ELS Method)
For every $x\in{\bf dom\,}\,g$ and $k\in\mathbbm{N}$, we have
$2\beta_{k}[F(x_{k})-F(x)]+2({F(x_{k+1})-F(x_{k})})\leq\|x_{k}-x\|^{2}-\|x_{k+1}-x\|^{2}-{\alpha\beta_{k}}\|x_{k}-\tilde{x}_{k}\|^{2}.$ (40)
Additionally, the sequence $(F(x_{k}))_{k\in\mathbbm{N}}$ is decreasing and convergent, and $\sum_{k=0}^{+\infty}\|x_{k}-x_{k+1}\|^{2}<+\infty$.
Proof Let $x\in{\bf dom\,}\,g$ and $k\in\mathbb{N}$. In view of the inexact relative inclusion (32), we have $v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k})\in\partial_{\varepsilon_{k}}g(\tilde{x}_{k})$, and hence the definition of $\partial_{\varepsilon}g$ in (5) implies that
$g(x)\geq g(\tilde{x}_{k})+\langle v_{k}+{x}_{k}-\tilde{x}_{k}-\nabla f({x}_{k}),x-\tilde{x}_{k}\rangle-\varepsilon_{k}.$ (41)
Since $f$ is convex, we have $f(x)-f(x_{k})\geq\langle\nabla f(x_{k}),x-x_{k}\rangle$. Adding the above two inequalities, using $f+g=F$, and simplifying the resulting expression, we obtain
$F(x)-F(x_{k})\geq g(\tilde{x}_{k})-g(x_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-x_{k}\rangle-\langle{{x}_{k}-\tilde{x}_{k}},{\tilde{x}_{k}-x}\rangle+\langle{v_{k}},{x-\tilde{x}_{k}}\rangle-\varepsilon_{k}.$
Combining the above inequality with the identity
$-\langle{{x}_{k}-\tilde{x}_{k}},{\tilde{x}_{k}-x}\rangle=\frac{1}{2}\left[\|x_{k}-\tilde{x}_{k}\|^{2}+\|\tilde{x}_{k}-x\|^{2}-\|x_{k}-x\|^{2}\right],$
we have
$F(x)-F(x_{k})\geq$ $\;g(\tilde{x}_{k})-g(x_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-x_{k}\rangle+\frac{1}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+\frac{1}{2}\left[\|\tilde{x}_{k}-x\|^{2}-\|x_{k}-x\|^{2}\right]+\langle{v_{k}},{x-\tilde{x}_{k}}\rangle-\varepsilon_{k},$
or, equivalently,
$F(x)-F(x_{k})+\frac{F(x_{k})-F(x_{k+1})}{\beta_{k}}\geq$ $\frac{g(x_{k})-g(x_{k+1})}{\beta_{k}}+g(\tilde{x}_{k})-g(x_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-y_{k}\rangle$
$+\frac{f(x_{k})-f(x_{k+1})}{\beta_{k}}+\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{1}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+\frac{1}{2}\left[\|x-(\tilde{x}_{k}-v_{k})\|^{2}-\|x_{k}-x\|^{2}\right]-\frac{1}{2}\|v_{k}\|^{2}-\varepsilon_{k}.$
Since $x_{k+1}=x_{k}+\beta_{k}(y_{k}-x_{k})$ and $g$ is convex, we have $g(x_{k+1})-g(x_{k})\leq\beta_{k}(g(y_{k})-g(x_{k}))$. Hence, combining the last two inequalities and the fact that $y_{k}=\tilde{x}_{k}-v_{k}$, we get
$F(x)-F(x_{k})+\frac{F(x_{k})-F(x_{k+1})}{\beta_{k}}\geq$ $\frac{1}{2}\left[\|x-y_{k}\|^{2}-\|x_{k}-x\|^{2}\right]$
$+\frac{f(x_{k})-f(x_{k+1})}{\beta_{k}}+\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{1}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+g(\tilde{x}_{k})-g(y_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-y_{k}\rangle-\frac{1}{2}\|v_{k}\|^{2}-\varepsilon_{k}.$
or, equivalently,
$F(x)-F(x_{k})$ $+\frac{F(x_{k})-F(x_{k+1})}{\beta_{k}}\geq\frac{1}{2}\left[\|x-y_{k}\|^{2}-\|x_{k}-x\|^{2}\right]+\frac{\alpha}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+\frac{f(x_{k})-f(x_{k+1})}{\beta_{k}}+\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{\tau}{2}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\|v_{k}\|^{2}+\gamma_{2}\varepsilon_{k}$
$+g(\tilde{x}_{k})-g(y_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-y_{k}\rangle-\frac{(1+\gamma_{1})}{2}\|v_{k}\|^{2}-(1+\gamma_{2})\varepsilon_{k}$
$+\frac{(1-\tau-\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2}.$
Now, using the linesearch procedure of Step 4, we obtain
$F(x)-F(x_{k})+\frac{F(x_{k})-F(x_{k+1})}{\beta_{k}}\geq$ $\frac{1}{2}\left[\|x-y_{k}\|^{2}-\|x_{k}-x\|^{2}\right]+\frac{\alpha}{2}\|x_{k}-\tilde{x}_{k}\|^{2}$
$+g(\tilde{x}_{k})-g(y_{k})+\langle\nabla f(x_{k}),\tilde{x}_{k}-y_{k}\rangle$
$-\frac{(1+\gamma_{1})}{2}\|v_{k}\|^{2}-(1+\gamma_{2})\varepsilon_{k}+\frac{(1-\tau-\alpha)}{2}\|x_{k}-\tilde{x}_{k}\|^{2}.$
It follows from the above inequality and (33) that
$\beta_{k}[F(x)-F(x_{k})]+{F(x_{k})-F(x_{k+1})}\geq\frac{\beta_{k}}{2}\left[\|x-y_{k}\|^{2}-\|x_{k}-x\|^{2}\right]+\frac{\alpha\beta_{k}}{2}\|x_{k}-\tilde{x}_{k}\|^{2}.$ (42)
On the other hand, using the identity $x_{k+1}-x=(1-\beta_{k})(x_{k}-x)+\beta_{k}(y_{k}-x)$ and the strong convexity of $\|\cdot\|^{2}$, we have
$\|x_{k+1}-x\|^{2}\leq(1-\beta_{k})\|x_{k}-x\|^{2}+\beta_{k}\|y_{k}-x\|^{2}-(1-\beta_{k})\beta_{k}\|x_{k}-y_{k}\|^{2},$
which implies
$\beta_{k}\left(\|y_{k}-x\|^{2}-\|x_{k}-x\|^{2}\right)\geq\|x_{k+1}-x\|^{2}-\|x_{k}-x\|^{2}.$
Therefore, the proof of (40) follows by combining the latter inequality with (42). The last statement of the proposition follows immediately from (40) with $x=x_{k}$ that
$02({F(x_{k+1})-F(x_{k})})\leq-\|x_{k+1}-x_{k}\|^{2}<$ 0 . (43)
So, the sequence $(F(x_{k}))_{k\in\mathbbm{N}}$ is decreasing and convergent because it is bounded from below by $F(x_{*})$. Moreover, the last inequality implies that $\sum_{k=0}^{+\infty}\|x_{k}-x_{k+1}\|^{2}<+\infty$. ∎
Next, we establish the full convergence of the sequence $(x_{k})_{k\in\mathbbm{N}}$ to a solution of problem (1). The proof is based on the quasi-Fejér convergence of the sequence $(x_{k})_{k\in\mathbbm{N}}$ to the set $S_{*}$, as defined in Definition 2. Note that to establish the convergence of $(x_{k})_{k\in\mathbbm{N}}$ we do not require any local or global Lipschitz continuity assumption of $\nabla f$.
Theorem 5.1 (Convergence for the IPG-ELS Method)
The sequence $(x_{k})_{k\in\mathbbm{N}}$ generated by the IPG-ELS method converges to a point in $S_{*}$.
Proof By employing Proposition 3 at $x=x_{*}\in S_{*}\subseteq{\bf dom\,}g$, we have
$\|x_{k+1}-x_{*}\|^{2}\leq\|x_{k}-x_{*}\|^{2}+2\left[F(x_{k})-F(x_{k+1})\right]\quad\mbox{for all}\quad k\in\mathbbm{N}.$ (44)
We set $\delta_{k}:=2\left[F(x_{k})-F(x_{k+1})\right]\geq 0$, and we will prove that $(\delta_{k})_{k\in\mathbbm{N}}$ is a summable sequence. In fact,
$\sum_{k=0}^{+\infty}\delta_{k}=$ $2\sum_{k=0}^{+\infty}\Big{[}F(x_{k})-F(x_{k+1})\Big{]}\leq 2\Big{[}F(x_{0})-\lim_{k\to+\infty}F(x_{k+1})\Big{]}$
$\leq$ $2\Big{[}F(x_{0})-F(x_{*})\Big{]}<+\infty.$
This together with (44) tells us that the sequence $(x_{k})_{k\in\mathbbm{N}}$ is quasi-Fejér convergent to $S_{*}$ via Definition 2. By Lemma 1(a), the sequence $(x_{k})_{k\in\mathbbm{N}}$ is bounded. Let $\bar{x}$ be an accumulation point of $(x_{k})_{k\in\mathbbm{N}}$. Hence, there exists a subsequence $(x_{\ell_{k}})_{k\in\mathbbm{N}}$ converging to $\bar{x}$.
Now we proceed by considering the two possible cases:
Case 1. The sequence $\left(\beta_{\ell_{k}}\right)_{k\in\mathbbm{N}}$ does not converge to 00, i.e., there exist some $\beta>0$ and a subsequence of $\left(\beta_{\ell_{k}}\right)_{k\in\mathbbm{N}}$ (without relabeling) such that
$\beta_{\ell_{k}}\geq\beta,\quad\forall\,k\in\mathbbm{N}.$ (45)
By using Proposition 3 with $x=x_{*}\in S_{*}$, we get
$\beta_{k}\left[F(x_{k})-F(x_{*})\right]\leq$ $\frac{1}{2}(\|x_{k}-x_{*}\|^{2}-\|x_{k+1}-x_{*}\|^{2})+F(x_{k})-F(x_{k+1}).$
Summing for $k=0,\ldots,m$, the above inequality implies that
$\sum_{k=0}^{m}\beta_{k}\left[F(x_{k})-F(x_{*})\right]\leq$ $\frac{1}{2}(\|x_{0}-x_{*}\|^{2}-\|x_{m+1}-x_{*}\|^{2})+F(x_{0})-F(x_{m+1})$
$\leq$ $\frac{1}{2}\|x_{0}-x_{*}\|^{2}+F(x_{0})-F(x_{*}).$
By taking $m\to+\infty$ and using the fact that $F(x_{k})\geq F(x_{*})$ and (45), we obtain that
$\beta\sum_{k=0}^{+\infty}\left[F(x_{\ell_{k}})-F(x_{*})\right]\leq\sum_{k=0}^{+\infty}\beta_{\ell_{k}}\left[F(x_{\ell_{k}})-F(x_{*})\right]\leq\sum_{k=0}^{+\infty}\beta_{k}\left[F(x_{k})-F(x_{*})\right]<+\infty,$
[…]
$F(x_{*})\leq F(\bar{x})\leq\liminf_{k\rightarrow+\infty}F(x_{\ell_{k}})=\lim_{k\rightarrow+\infty}F(x_{\ell_{k}})=F(x_{*}),$
which yields $F(\bar{x})=F(x_{*})$ and thus $\bar{x}\in S_{*}$.
[…]
$\hat{x}_{k+1}:=x_{k}+\hat{\beta}_{k}(y_{k}-x_{k})=(1-\hat{\beta}_{k})x_{k}+\hat{\beta}_{k}y_{k},$ (46)
where $y_{k}=\tilde{x}_{k}-v_{k}$. It follows from the definition of the linesearch that
$f(\hat{x}_{k+1})>f(x_{k})+\hat{\beta}_{k}\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\frac{\tau}{2}\hat{\beta}_{k}\|x_{k}-\tilde{x}_{k}\|^{2}+\frac{\gamma_{1}}{2}\hat{\beta}_{k}\|v_{k}\|^{2}+\hat{\beta}_{k}\gamma_{2}\varepsilon_{k}.$ (47)
It follows from the convexity of $f$, the fact that $\gamma_{1}>1>\tau$, and the positiveness of the term $\hat{\beta}_{k}\gamma_{2}\varepsilon_{k}$ that
$\langle\nabla f(\hat{x}_{k+1}),\hat{x}_{k+1}-x_{k}\rangle\geq f(\hat{x}_{k+1})-f(x_{k})>\hat{\beta}_{k}\langle\nabla f(x_{k}),y_{k}-x_{k}\rangle+\hat{\beta}_{k}\frac{\tau}{2}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)$
which, together with (46), yields
$\hat{\beta}_{k}\frac{\tau}{2}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)$ $<\hat{\beta}_{k}\langle\nabla f(\hat{x}_{k+1})-\nabla f(x_{k}),y_{k}-x_{k}\rangle$
$\leq\hat{\beta}_{k}\|\nabla f(\hat{x}_{k+1})-\nabla f(x_{k})\|\|y_{k}-x_{k}\|$
$=\hat{\beta}_{k}\|\nabla f(\hat{x}_{k+1})-\nabla f(x_{k})\|\|x_{k}-\tilde{x}_{k}+v_{k}\|.$
Now it follows from Lemma 2 that
$\|x_{k}-\tilde{x}_{k}+v_{k}\|^{2}\leq 2\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right).$
Hence,
$\hat{\beta}_{k}\frac{\tau}{2}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)<\hat{\beta}_{k}\|\nabla f(\hat{x}_{k+1})-\nabla f(x_{k})\|\cdot\sqrt{2}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)^{\frac{1}{2}},$
which, due to the positiveness of $\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}$, yields
$\frac{\tau\sqrt{2}}{4}\left(\|x_{k}-\tilde{x}_{k}\|^{2}+\|v_{k}\|^{2}\right)^{\frac{1}{2}}\leq\|\nabla f(\hat{x}_{k+1})-\nabla f(x_{k})\|.$ (48)
Note that $\|\hat{x}_{k+1}-x_{k}\|=\|\hat{\beta}_{k}(y_{k}-x_{k})\|=\frac{\hat{\beta}_{k}}{\beta_{k}}\|x_{k+1}-x_{k}\|=\frac{1}{\theta}\|x_{k+1}-x_{k}\|,$ which combined with the last statement of Proposition 3 give us that $\|\hat{x}_{\ell_{k}+1}-x_{\ell_{k}}\|\to 0$ as $k\to+\infty$. Since $\nabla f$ is continuous, we have $\|\nabla f(\hat{x}_{\ell_{k}+1})-\nabla f(x_{\ell_{k}})\|\to 0$ as $k\to+\infty$. From (48), it is derived that
$\lim_{k\rightarrow+\infty}\,\|x_{\ell_{k}}-\tilde{x}_{\ell_{k}}\|^{2}+\|v_{\ell_{k}}\|^{2}\leq 0,$ (49)
therefore, $\lim_{k\rightarrow+\infty}\,\|x_{\ell_{k}}-\tilde{x}_{\ell_{k}}\|=0$ and $\lim_{k\rightarrow+\infty}\|v_{\ell_{k}}\|=0$. Additionally, we can use (35) to show that $\lim_{k\rightarrow+\infty}\varepsilon_{\ell_{k}}=0$. Thus, $\bar{x}$ is also an accumulation point of the sequence $(\tilde{x}_{k})_{k\in\mathbb{N}}$, and $\tilde{x}_{\ell_{k}}\to\bar{x}$ as $k\to+\infty$. Moreover, we have that
$\lim_{k\rightarrow+\infty}\|\nabla f(x_{\ell_{k}})-\nabla f(\tilde{x}_{\ell_{k}})\|=0.$ (50)
Now, using (33), we obtain
$w_{k}:=x_{k}-\tilde{x}_{k}+v_{k}+\nabla f(\tilde{x}_{k})-\nabla f(x_{k})\in\nabla f(\tilde{x}_{k})+\partial_{\varepsilon_{k}}g(\tilde{x}_{k})\subseteq\partial_{\varepsilon_{k}}F(\tilde{x}_{k}).$
Furthermore, since $v_{\ell_{k}}$ converges to 00 as indicated by (49), and by applying the triangular inequality, we obtain
$\|w_{k}\|=\|x_{k}-\tilde{x}_{k}+v_{k}+\nabla f(\tilde{x}_{k})-\nabla f(x_{k})\|\leq\|x_{k}-\tilde{x}_{k}\|+\|v_{k}\|+\|\nabla f(\tilde{x}_{k})-\nabla f(x_{k})\|,$
which implies, via (49) and (50), that $w_{\ell_{k}}\in\partial_{\varepsilon_{\ell_{k}}}F(\tilde{x}_{\ell_{k}})$ also converges to 00. Consequently, the convergence of $\tilde{x}_{\ell_{k}}$ to $\bar{x}$ and $\varepsilon_{\ell_{k}}$ to 00, combined with the closedness of the graph of $\partial F$ in Proposition 1, gives us that $0\in\partial F(\bar{x})$. This is equivalent to stating that $\bar{x}\in S_{*}$.
In all the cases considered above, we have shown that $\bar{x}$, an accumulation point of the sequence $(x_{k})_{k\in\mathbb{N}}$, belongs to $S_{*}$. Proposition 1(b) implies that $(x_{k})_{k\in\mathbb{N}}$ converges to an optimal solution in $S_{*}$. ∎
We continue by showing the convergence rate of the functional values sequence $(F(x_{k}))_{k\in\mathbbm{N}}$.
Theorem 5.2 (Convergence Rate of the IPG-ELS Method)
Let $(x_{k})_{k\in\mathbbm{N}}$ and $(\beta_{k})_{k\in\mathbbm{N}}$ be generated by the IPG-ELS method. Assume that there exists $\beta>0$ such that $\beta_{k}\geq\beta$ for all $k\in\mathbbm{N}$. Then, for all $k\in\mathbbm{N}$, we have
$F(x_{k})-F_{*}\leq\frac{{\rm dist}(x_{0},S_{*})^{2}+2\left(F(x_{0})-F_{*}\right)}{2\beta(k+1)}.$ (51)
Proof For any $\ell\in\mathbbm{N}$ and $x_{*}\in S_{*}$, it follows from Proposition 3 with $k=\ell$ and $x=x_{*}$ that
$F(x_{\ell})-F(x_{*})\leq\frac{1}{2\beta_{\ell}}\left[\|x_{\ell}-x_{*}\|^{2}-\|x_{\ell+1}-x_{*}\|^{2}+2\left(F(x_{\ell})-F(x_{\ell+1})\right)\right]$
Summing the above inequality over $\ell=0,1,\ldots,k$, we have
$\sum_{\ell=0}^{k}\left(F(x_{\ell})-F(x_{*})\right)$ $\leq\frac{1}{2\beta}\left[\|x_{0}-x_{*}\|^{2}-\|x_{k+1}-x_{*}\|^{2}+2\left(F(x_{0})-F(x_{k+1})\right)\right]$
$\leq\frac{1}{2\beta}\left[\|x_{0}-x_{*}\|^{2}+2\left(F(x_{0})-F(x_{*})\right)\right].$ (52)
Since in view of the last statement of Proposition 3, we have $F(x_{\ell+1})\leq F(x_{\ell})$ for any $\ell\in\mathbbm{N}$, it follows from (5) that
$(k+1)\left(F(x_{k})-F_{*}\right)\leq\frac{1}{2\beta}\left(\|x_{0}-x_{*}\|^{2}+2\left[F(x_{0})-F_{*}\right]\right).$
Since $x_{*}\in S_{*}$ is arbitrary, the proof of (51) follows. ∎
Remark 3 (Complexity of $\eta$-Approximate Solution)
It follows from Theorem 5.2 that, given any $\eta>0$, the IPG-ELS method generates an $\eta$-approximate solution $x_{k}$ to problem (1), in the sense that $F(x_{k})-F_{*}\leq\eta$ in at most $k=\mathcal{O}(1/\eta)$ iterations. We note further that Lemma 7 guarantees that the locally Lipschitz assumption for the gradient of $f$ can be used as a sufficient condition for establishing the convergence rate in Theorem 5.2.
We end this section by proving the complexity of an $\eta$-approximate stationary solution for problem (1) as in Definition 1. For this complexity result, we assume that $\nabla f$ is $L$-Lipschitz continuous on the domain of $g$.
Theorem 5.3 (Complexity of $\eta$-Approximate Stationary Solution)
Consider $(x_{k})_{k\in\mathbbm{N}}$, $(\tilde{x}_{k})_{k\in\mathbbm{N}}$, $(v_{k})_{k\in\mathbbm{N}}$, and $(\varepsilon_{k})_{k\in\mathbbm{N}}$ generated by the IPG-ELS method and define $w_{k}:=v_{k}+x_{k}-\tilde{x}_{k}+\nabla f(\tilde{x}_{k})-\nabla f(x_{k})$, for every $k\in\mathbbm{N}$. Then, we have
$w_{k}\in\nabla f(\tilde{x}_{k})+\partial_{\varepsilon_{k}}g(\tilde{x}_{k})\subseteq\partial_{\varepsilon_{k}}F(\tilde{x}_{k}),\qquad\forall k\in\mathbbm{N}.$ (53)
Additionally, if $\alpha\in(0,1-\tau)$, and $\nabla f$ is $L$-Lipschitz continuous on ${\bf dom\,}g$, then, given a tolerance $\eta>0$, the IPG-ELS method generates an $\eta$-approximate stationary solution $\tilde{x}_{k}$ to problem (1) with residues $(w_{k},\varepsilon_{k})$, in the sense of Definition 1, in at most $k=O\left(1/\eta^{2}\right)$ iterations.
Proof The first inclusion in (53) follows immediately from (32) and the definition of $w_{k}$, whereas the second inclusion in (53) follows from the definitions of $F$ and $\partial_{\varepsilon}F$.
Now let $x_{*}$ be the projection of $x_{0}$ onto $S_{*}$ and let $d_{0}:=\|x_{0}-x_{*}\|$. As it was observed in Lemma 6, the Lipschitz continuity of $\nabla f$ implies that there exists $\beta>0$ such that $\beta_{k}\geq\beta$ for all $k$. It follows from Proposition 3 with $k=\ell\in\mathbbm{N}$ and $x=x_{*}\in S_{*}$ that
$\alpha\beta\|x_{\ell}-\tilde{x}_{\ell}\|^{2}\leq\left(\|x_{\ell}-x_{*}\|^{2}-\|x_{\ell+1}-x_{*}\|^{2}+2\left[F(x_{\ell})-F(x_{\ell+1})\right]\right)$ (54)
for all $\ell\in\mathbbm{N}$. Summing the above inequality over $\ell=0,1,\ldots,k$, and using that $x_{*}\in S_{*}$, we have
$\alpha\beta\sum_{\ell=0}^{k}\|x_{\ell}-\tilde{x}_{\ell}\|^{2}\leq\|x_{0}-x_{*}\|^{2}+2\left[F(x_{0})-F(x_{*})\right].$
Hence, since $d_{0}=\|x_{0}-x_{*}\|$, we see that there exists ${\ell_{k}}\leq k$ such that
$\|x_{{\ell_{k}}}-\tilde{x}_{{\ell_{k}}}\|^{2}\leq\frac{d_{0}^{2}+2[F(x_{0})-F(x_{*})]}{\alpha\beta(k+1)}.$ (55)
On the other hand, since $\tau,\alpha>0$, if follows from (35) that, for every $\ell\in\mathbbm{N}$,
$\frac{(\gamma_{1}-1)}{2}\|v_{\ell}\|^{2}+\gamma_{2}\varepsilon_{\ell}$ $\leq\frac{\|x_{\ell}-\tilde{x}_{\ell}\|^{2}}{2}+\langle x_{\ell}-\tilde{x}_{\ell},v_{\ell}\rangle$
$\leq\frac{\|x_{\ell}-\tilde{x}_{\ell}\|^{2}}{2}+\frac{\|x_{\ell}-\tilde{x}_{\ell}\|^{2}}{\gamma_{1}-1}+\frac{(\gamma_{1}-1)}{4}\|v_{\ell}\|^{2},$
where the last inequality is due to Cauchy-Schwarz inequality and the fact that $ab\leq sa^{2}/2+b^{2}/(2s)$ for any $a,b\in\mathbbm{R}$ and $s>0$, in particular, with $a=\|v_{\ell}\|$, $b=\|\tilde{x}_{\ell}-x_{\ell}\|$, and $s=(\gamma_{1}-1)/2$. Hence, we have
$\frac{(\gamma_{1}-1)}{4}\|v_{\ell}\|^{2}+\gamma_{2}\varepsilon_{\ell}\leq\left(\frac{\gamma_{1}+1}{2(\gamma_{1}-1)}\right)\|\tilde{x}_{\ell}-x_{\ell}\|^{2},\qquad\forall\ell\in\mathbbm{N}.$ (56)
Now, from the definition of $w_{\ell}$, (56), the Cauchy-Schwarz inequality, the fact that $x_{k},\tilde{x}_{k}$ are in the domain of $g$, and the $L$-Lipschitz continuity of $\nabla f$ on the domain of $g$, we have, for every $\ell\in\mathbbm{N}$,
$\|w_{\ell}\|$ $\leq\|v_{\ell}\|+\|{x}_{\ell}-\tilde{x}_{\ell}\|+\|\nabla f(x_{\ell})-\nabla f(\tilde{x}_{\ell})\|$
$\leq\left[\frac{\sqrt{2(\gamma_{1}+1)}}{\gamma_{1}-1}+1+L\right]\|x_{\ell}-\tilde{x}_{\ell}\|.$ (57)
Moreover, it follows from (56) that
$\varepsilon_{\ell}\leq\left(\frac{\gamma_{1}+1}{2\gamma_{2}(\gamma_{1}-1)}\right)\|x_{\ell}-\tilde{x}_{\ell}\|^{2}.$ (58)
Hence, it follows from (55), (5), and (58) with $\ell=\ell_{k}$ and $m_{0}:=d_{0}^{2}+2(F(x_{0})-F(x_{*}))$ that
$w_{\ell_{k}}\leq\left[\frac{\sqrt{2(\gamma_{1}+1)}}{\gamma_{1}-1}+1+L\right]\frac{\sqrt{m_{0}}}{\sqrt{\alpha\beta(k+1)}},\qquad\varepsilon_{\ell_{k}}\leq\left(\frac{\gamma_{1}+1}{2\gamma_{2}(\gamma_{1}-1)}\right)\frac{m_{0}}{\alpha\beta(k+1)}$
which in turn implies that
$w_{\ell_{k}}=\mathcal{O}(1/\sqrt{k}),\quad\varepsilon_{\ell_{k}}=\mathcal{O}(1/k).$ (59)
Thus, the last statement of the theorem follows from (59) and the first inclusion in (53). ∎
6 Numerical Experiments
In this section, we investigate the numerical behavior of the IPG-ELS method in solving the CUR-like factorization optimization problem JMLR:v12:mairal11a . Consider $\mathbbm{E}=\mathbbm{R}^{n\times m}$. Given a matrix $W\in\mathbbm{R}^{m\times n}$, the objective is to find a matrix with sparse rows and columns, $X\in\mathbbm{R}^{n\times m}$, such that $WXW$ approximates $W$. This problem can be formulated as the following splitting optimization problem:
$\min_{X\in\mathbbm{R}^{n\times m}}\left\{\frac{1}{2}\|W-WXW\|_{F}^{2}+\lambda_{\text{row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+\lambda_{\text{col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2}\right\},$ (60)
where $\|\cdot\|_{F}$ denotes the Frobenius norm, and $X^{(i)}$ and $X_{(j)}$ denote the $i$-th row and $j$-th column of $X$, respectively. This problem is a special case of problem (1) with
$f(X):=\frac{1}{2}\|W-WXW\|_{F}^{2},\quad g(X):=\lambda_{\text{row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+\lambda_{\text{col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2}.$
In this case, the gradient of $f$ is given by $\nabla f(X)=W^{T}(WXW-W)W^{T}$ and has a Lipschitz constant $L=\|W^{T}W\|_{F}^{2}$. Note that the proximal operator of $g$ does not have a closed-form solution; however, $g$ exhibits a separable structure. By following the approach described in Section 3.1, we compute a triple $(\tilde{X}_{k},0,\varepsilon_{k})$ that satisfies conditions (32)-(33) using the Dykstra-like algorithm (Bauschke:2008c, , Theorem 3.3). This algorithm is applied to the proximal subproblem
$\min_{X\in\mathbbm{R}^{n\times m}}\left\{\frac{1}{2}\|X-Z\|^{2}_{F}+{\lambda_{\rm row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+{\lambda_{\rm col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2}\right\},$ (61)
where $z:=X_{k}-\nabla f(X_{k}),$ with the initial points $z_{0}=z,\quad p_{0}=0,\quad q_{0}=0$, generates the sequences
$\left\{\begin{array}[]{l}y_{\ell}={\bf prox}_{g_{1}}(z_{\ell}+p_{\ell})\\ p_{\ell+1}=z_{\ell}+p_{\ell}-y_{\ell}\end{array}\right.\quad\mbox{and}\quad\left\{\begin{array}[]{l}z_{\ell+1}={\bf prox}_{g_{2}}(y_{\ell}+q_{\ell})\\ q_{\ell+1}=y_{\ell}+q_{\ell}-z_{\ell+1},\end{array}\right.$
where $g_{1}(x)=\lambda_{\rm col}\sum_{j=1}^{m}\|x_{(j)}\|_{2}$ and $g_{2}(x)=\lambda_{\rm row}\sum_{i=1}^{n}\|x^{(i)}\|_{2}$. It follows from the separable property of the proximal operator and Section 6.5.1 of Parikh-Boyd:2014 that
${\bf prox}_{g_{1}}(x)_{(j)}=\max\left\{1-\frac{\lambda_{\text{col}}}{\|x_{(j)}\|_{2}},0\right\}x_{(j)},\;\text{for }\quad j=1,\ldots,m$
and
${\bf prox}_{g_{2}}(x)^{(i)}=\max\left\{1-\frac{\lambda_{\text{row}}}{\|x^{(i)}\|_{2}},0\right\}x^{(i)},\;\text{for }\quad i=1,\ldots,n.$
Hence, if the Dykstra-like algorithm is terminated when
$\varepsilon_{\ell}\leq\frac{(1+\gamma_{2})(1-\tau-\alpha)}{2}\|X_{k}-z_{\ell}\|^{2},$
then $\tilde{X}_{k}:=z_{\ell}$ is an inexact proximal solution of (61) with residual $(V_{k},\varepsilon_{k}):=(0,\varepsilon_{\ell})$, where $\varepsilon_{\ell}$ is as defined in (17).
Considering that the IPG-ELS method integrates two effective strategies: (i) permitting the subproblem to be solved inexactly to meet a relative approximation criterion, and (ii) employing an explicit linesearch procedure that computes the proximal operator only once per iteration, our primary goal is to demonstrate that, in certain cases, the proposed method surpasses the proximal gradient method that employs only one of these strategies. Consequently, we compare the new algorithm with three alternative schemes: an exact version of the IPG-ELS method, denoted by PG-ELS, which corresponds to IPG-ELS with $\gamma_{1}=\gamma_{2}=0$, $\theta=0.5$, $\tau=1$, and $\epsilon_{k}\ \leq 10^{-12}$, replacing the inexact criterion in (33) with $v_{k}=0$; an IPG method with a fixed Stepsize, corresponding to (Millan:2019, , Algorithm 2) with $\alpha_{k}=1/L$, $\bar{w}_{k}=L(X_{k}-\tilde{X}_{k})-\nabla f(X_{k})$ and $\sigma^{2}=0.9$, where $L$ is the Lipschitz constant of $f$. This algorithm is denoted by IPG-FixStep and is defined as $X_{k+1}=\tilde{X}_{k},$ for all $k\geq 0,$ where $\tilde{X}_{k}$ satisfies
$0\in\partial_{\varepsilon_{k}}g(\tilde{X}_{k})+L(\tilde{X}_{k}-X_{k})+\nabla f(X_{k}),\quad{\varepsilon_{k}}\leq{0.9L}\|\tilde{X}_{k}-\left(X_{k}-(1/L)\nabla f(x_{k})\right)\|^{2}/2;$
and an instance of Tseng’s modified forward-backward splitting method, as described in Monteiro:2010 , applied to (60). The stepsize is fixed at $0.9/L$, and the triple $(\tilde{X}_{k},0,\varepsilon_{k})$, associated with the proximal subproblem, is computed to ensure $\varepsilon_{k}\leq 10^{-12}$. This algorithm will be referred to as Tseng-MFBS.
For all methods, the approximate proximal solution $\tilde{X}_{k}$ of (61), along with its residual $(0,\varepsilon_{k})$, is computed using the Dykstra-like algorithm as described above. The initialization parameters for the IPG-ELS method were set as $\tau=0.8$, $\theta=0.5$, $\gamma_{1}=\gamma_{2}=1.1$, and $\alpha=0.01$. For all tests, we initialized $X_{0}=0\in\mathbbm{R}^{n\times m}$, and set $\lambda_{\text{row}}=\lambda_{\text{col}}=0.01$. The IPG-ELS method was executed for $101$ outer iterations to establish a baseline performance metric, $F_{*}:=F(X_{101})$. The other three algorithms were terminated as soon as $F(X_{k})\leq F_{*}$ or after reaching a maximum of 2001 iterations.. The algorithms were evaluated on six datasets from Cano:2005 ; Guyon:2004 ; Team:2008 : Colon tumor ($62\times 2000$), heart disease ($303\times 14$), central nervous system (CNS) ($60\times 7129$), lung cancer-Michigan ($96\times 7129$), Secom ($1567\times 590$), and Cina0 ($132\times 16033$).
Each matrix $W$ was normalized to have a unit Frobenius norm, with an additional step of centering each column. Subsequently, the resulting matrices were multiplied by a constant, which plays a crucial role in controlling the Lipschitz constant of the function $f$. The experiments were conducted using the Python programming language, which was installed on a machine equipped with a 3.5 GHz Dual-Core Intel Core i5 processor and 16 GB of 2400 MHz DDR4 memory.
In Tables 1 and 2, we report the Lipschitz constant of the gradient of $f$ (denoted as Lips), the number of outer iterations (O-IT), the number of inner iterations (I-IT), the number of linesearch iterations (LS-IT), and the total running time in seconds (Time). The results indicate that, in terms of CPU times, the IPG-ELS method outperforms the other three methods. This efficiency can be attributed to two main factors: (i) the PG-ELS and Tseng-MFBS methods require significantly more inner iterations to solve the proximal subproblem ”exactly”, and (ii) the IPG-FixStep and Tseng-MFBS methods employ small stepsizes of $1/L$ and $0.9/L$, respectively, in the gradient step.
Problem Lips Method $F(X_{k})$ O-IT I-IT LS-IT Time
Colon Tumor 41.58 IPG-ELS 1.1056 101 195 318 25.80
PG-ELS 1.1056 104 751 324 32.82
IPG-FixStep 1.1056 1450 1451 - 137.78
Tseng-MFBS 1.1056 1615 4845 - 193.92
Colon Tumor 665.32 IPG-ELS 2.3647 101 101 823 31.02
PG-ELS 2.3623 128 455 1038 43.40
IPG-FixStep 2.3647 1103 1104 - 94.51
Tseng-MFBS 2.3645 1227 2454 - 116.80
Colon Tumor 5133.69 IPG-ELS 5.8989 101 101 1134 49.30
PG-ELS 5.8832 119 328 1324 63.73
IPG-FixStep 5.8981 586 587 - 66.26
Tseng-MFBS 5.8980 652 1304 - 79.46
Heart Disease 77.12 IPG-ELS 0.1732 101 178 539 1.25
PG-ELS 0.1732 119 876 603 3.68
IPG-FixStep 0.1732 487 488 - 2.92
Tseng-MFBS 0.1732 541 1082 - 4.79
Heart Disease 1233.99 IPG-ELS 0.3129 101 101 903 0.86
PG-ELS 0.3119 89 368 788 1.71
IPG-FixStep 0.3520 2001 2001 - 11.87
Tseng-MFBS 0.3577 2001 4002 - 17.56
Heart Disease 9521.56 IPG-ELS 0.4995 101 101 1222 1.01
PG-ELS 0.4992 227 622 2690 3.80
IPG-FixStep 0.7242 2001 2001 - 14.85
Tseng-MFBS 0.7533 2001 4002 - 20.74
CNS 41.95 IPG-ELS 0.9519 101 182 341 397.09
PG-ELS 0.9518 119 714 396 531.26
IPG-FixStep 0.9519 1217 1218 - 1762.69
Tseng-MFBS 0.9519 1354 3091 - 2263.33
CNS 671.17 IPG-ELS 2.1153 101 101 768 554.27
PG-ELS 2.1119 148 725 1110 885.04
IPG-FixStep 2.2193 2001 2001 - 2752.47
Tseng-MFBS 2.2439 2001 4002 - 2828.31
CNS 5178.78 IPG-ELS 6.0665 101 101 1130 854.13
PG-ELS 6.0613 102 343 1138 896.40
IPG-FixStep 6.0655 775 776 - 1240.28
Tseng-MFBS 6.0653 862 1724 - 1430.63
Table 1: Performance of the IPG-ELS, PG-ELS, IPG-FixStep and Tseng-MFBS methods on $3$ data sets.
Problem Lips Method $F(X_{k})$ O-IT I-IT LS-IT Time
Lung cancer 52.58 IPG-ELS 0.8985 101 179 443 539.37
PG-ELS 0.8984 105 837 457 646.59
IPG-FixStep 0.8985 678 679 - 1216.98
Tseng-MFBS 0.8985 755 2265 - 1487.95
Lung cancer 841.23 IPG-ELS 2.7632 101 101 845 601.68
PG-ELS 2.7623 131 768 1085 875.26
IPG-FixStep 2.7631 588 589 - 850.73
Tseng-MFBS 2.7631 654 1308 - 978.97
Lung cancer 2658.70 IPG-ELS 3.6391 101 101 992 711.25
PG-ELS 3.6378 161 740 1574 1398.46
IPG-FixStep 3.8819 2001 2001 - 3437.99
Tseng-MFBS 3.9389 2001 4002 - 3124.96
Secom 45.78 IPG-ELS 0.6438 101 175 373 99.15
PG-ELS 0.6438 99 9247 360 822.17
IPG-FixStep 0.6438 857 858 - 304.24
Tseng-MFBS 0.6438 952 34344 - 5216.87
Secom 732.51 IPG-ELS 0.8587 101 101 779 86.11
PG-ELS 0.8586 102 4931 795 431.53
IPG-FixStep 0.8586 1662 1663 - 454.59
Tseng-MFBS 0.8587 1847 3969 - 637.36
Secom 5652.07 IPG-ELS 1.6981 101 101 1138 108.07
PG-ELS 1.6899 125 1346 1388 220.10
IPG-FixStep 1.6977 801 802 - 218.15
Tseng-MFBS 1.6977 891 1782 - 296.50
Cina0 68.39 IPG-ELS 0.7972 101 245 487 3168.99
PG-ELS 0.7972 104 1490 483 5925.88
IPG-FixStep 0.7972 609 610 - 6451.18
Tseng-MFBS 0.7972 677 1354 - 8000.69
Cina0 527.70 IPG-ELS 1.2817 101 250 838 4567.13
PG-ELS 1.2817 94 1789 693 5704.39
IPG-FixStep 1.2817 1010 1011 - 8735.91
Tseng-MFBS 1.2817 1122 2244 - 11564.76
Cina0 8443.20 IPG-ELS 3.6531 101 104 1126 3968.04
PG-ELS 3.6530 284 3083 3061 15889.08
IPG-FixStep 3.8493 2001 2001 - 17796.07
Tseng-MFBS 3.8657 2001 4002 - 20681.39
Table 2: Performance of the IPG-ELS, PG-ELS, IPG-FixStep and Tseng-MFBS methods on $3$ data sets.
7 Conclusions
In this work, we present an inexact proximal gradient method for solving composite convex optimization problems. This method features a novel explicit linesearch using the relative-type inexact solution of the proximal subproblem. Our approach is primarily designed to solve splitting problems when the objective function is the sum of differentiable and nondifferentiable convex functions, and the analytical computation of the proximal operator is not available. Notably, the convergence of the proposed method is established without assuming Lipschitz continuity of the gradient of the smooth function. This method addresses the need for a balance between computational efficiency and the accuracy of solving the proximal subproblem, a common challenge in practice.
We have confirmed the convergence and iteration complexity of our method, validating its theoretical soundness and practical utility. Numerical experiments demonstrate its applicability and efficiency. Our method maintains convergence rates while efficiently managing relative inexact solutions of the proximal operator. The numerical results indicate that the proposed method competes effectively with both the exact proximal gradient method and the inexact proximal gradient method with a fixed stepsize.
Declarations
Funding
YBC
was partially supported by the NSF Grant DMS-2307328, and by an internal grant from NIU. MLNG was partially supported by CNPq Grants 405349/2021-1 and 304133/2021-3. JGM was partially supported by CNPq Grant 312223/2022-6.
Data availability statement
The codes supporting the numerical experiments are freely available in the homepage:
https://maxlng.ime.ufg.br/p/17888-publications.
Conflicts of interest
The authors declare that they have no conflict of interest.
Acknowledgements.
The authors are grateful to the anonymous referees for their valuable comments and suggestions which helped to improve the quality of the paper.